### v1

# Sparse-DeRF: Deblurred Neural Radiance Fields from Sparse View

Abstract
Recent studies construct deblurred neural radiance fields (DeRF) using dozens of blurry images, which are not practical scenarios if only a limited number of blurry images are available. This paper focuses on constructing DeRF from sparse-view for more pragmatic real-world scenarios. As observed in our experiments, establishing DeRF from sparse views proves to be a more challenging problem due to the inherent complexity arising from the simultaneous optimization of blur kernels and NeRF from sparse view. Sparse-DeRF successfully regularizes the complicated joint optimization, presenting alleviated overfitting artifacts and enhanced quality on radiance fields. The regularization consists of three key components: Surface smoothness, helps the model accurately predict the scene structure utilizing unseen and additional hidden rays derived from the blur kernel based on statistical tendencies of real-world; Modulated gradient scaling, helps the model adjust the amount of the backpropagated gradient according to the arrangements of scene objects; Perceptual distillation improves the perceptual quality by overcoming the ill-posed multi-view inconsistency of image deblurring and distilling the pre-filtered information, compensating for the lack of clean information in blurry images. We demonstrate the effectiveness of the Sparse-DeRF with extensive quantitative and qualitative experimental results by training DeRF from 2-view, 4-view, and 6-view blurry images.
Index Terms:
Neural Radiance Fields, Deblurring, Novel View Synthesis, 3D Synthesis, Neural Rendering, Sparse View setting
I Introduction
Representing 3-dimensional (3D) space from multi-view images has rapidly grown after the emergence of the neural radiance fields (NeRF), which maps continuous spatial coordinates to volume density and radiance fields. Its realistic rendering quality and simple architecture have led to widespread applications and collaborations with various research fields in computer vision and graphics. As practical applications of NeRF continue to attract attention, research in real-world scenarios has emerged as a promising research direction such as NeRF from noisy images or sparse view.
In real-world scenarios, tackling the blurry images from camera motion is regarded to be important since users often encounter degraded images when capturing photos with their own devices due to the unintentional camera movement during exposure time. To solve this problem, several NeRF studies [1, 2, 3] have attempted to construct deblurred neural radiance fields (hereafter, DeRF) from blurry images using joint optimization of internal implicit blur kernel and radiance fields, but they use dozens of blurry images to train, which is actually not practical scenarios. The assumed experimental environments, where radiance fields are trained from about 20 to 30 blurry images, seem unlikely to occur in reality Hence we delve into the practical consideration for situations where only blurry images are utilized. We reasoned that situations requiring the use of only blurry images would arise when the available images for reconstructing the desired 3D space are both very limited and blurry. Following this rationale, we propose a novel pragmatic scenario for radiance fields from blurry images that establish the DeRF from sparse view settings. Specifically, we set the 2-view, 4-view, and 6-view settings based on our consideration of the practical applications of research on generating radiance fields from blurry images.
Actually, the NeRF system already has an inherent drawback: it is prone to be overfitted to training views and struggles to grasp correct geometry when only sparse view inputs are available. Moreover, we experimentally find that blurred images lead to more severe overfitting in DeRF from sparse view because blur kernels introduce a more complex optimization process compared to standard NeRF. Due to the increased complexity, DeRF training suffers more structural distortion than general NeRFs when trained from sparse view, exhibiting further overfitting with floating artifacts as shown in our experiments. Although there are several works [4, 5, 6] to regularize radiance fields in sparse view scenarios, existing regularization methods are not effective in addressing the complex optimization issue of DeRF as demonstrated in comparative experiments using existing representative regularization techniques in sparse view NeRF and blur kernel of the DeRF, namely RegNeRF [6] and DP-NeRF [2]. Furthermore, in the DeRF system, it is challenging to use other data-deriven priors such as predicted depth supervision since available images do not ensure the confidence of the estimated values due to the inherent degradation of the given images. Therefore, our goal is to regularize the complex joint optimization of blur kernel and radiance fields for DeRF to enhance the structural and perceptual quality of radiance fields from sparse blurry images, overcoming the aforementioned challenging issues.
In this paper, we propose for the first time to ameliorate the spatial ambiguity and enhance the sharp texture of the DeRF from sparse view, which we refer to as Sparse-DeRF. We introduce a novel regularization method for easing complex joint optimization, which consists of two geometric constraints and a perceptual prior. Geometric constraints are proposed to predict the accurate structure in radiance fields from sparse view, which consists of surface smoothness (SS) and modulated gradient scaling (MGS). First, SS rectifies the overall geometry based on classical depth smoothness on integrated unobserved rays as similar to RegNeRF [6]. We utilize the novel hidden rays in camera motion cues derived from blur kernels as additional out-of-distribution unobserved rays to reflect the statistical flatness of real-world geometry as [7, 6] argued. Second, MGS is designed to flexibly modulate the scaling function to compensate for the gradients based on the arrangement of the scene components, which cannot be handled by a single scaling function in non-parameterized coordinate systems such as normalized device coordinates (NDC). It alleviates the spatial ambiguity arising from ray sampling and the disproportionate gradients of NeRF by introducing a parameterized sinusoidal function as a novel scaling function. These two geometric constraints improve the structural scene geometry of radiance fields even without explicit depth supervision in a sparse view setting.
In addition to geometric constraints, we propose the perceptual distillation (PD) as a perceptual prior to enhance the detailed texture of the radiance fields by taking advantage of the previously established image deblurring algorithm. Traditional image deblurring has shown significant performance improvements alongside the advancement of deep learning, demonstrating more enhanced details and textures. We believe that the sharp texture information from such deblurred images can be used as additional complementary information to achieve high fidelity in the Sparse-DeRF environment, where only a few degraded images are available to reconstruct the scene. However, while we can take the pre-filtered images with a pre-trained deep learning-based image deblurring model, the independence of image deblurring poses challenges in directly utilizing deblurred images as pixel-wise color supervision, due to inconsistency across the given images. This inconsistency comes from the inherent ill-posed property of the image deblurring that breaks the geometric and appearance consistency across the multi-view images of the single 3D scene. Hence, we impart the perceptual information of pre-filtered images to the radiance fields by distilling the features extracted from the deep learning-based image feature extractor. Extracted features enable the radiance fields to enhance perceptual quality by utilizing pre-deblurred textures.
Our results illustrate that the Sparse-DeRF produces high-quality rendered images from sparse blurry images, with improved perceptual texture quality and well-structured scene geometry. Additionally, we demonstrate the effectiveness of the proposed constraints and a prior through experimental results and analysis. Furthermore, we conduct comprehensive experiments to investigate ablations using two types of representative blur kernels from Deblur-NeRF [1] and DP-NeRF [2]. These experiments aim to show the superiority of the proposed regularization method and analyze its effects depending on the type of kernel employed.
II Related Work
II-A Neural Rendering and Radiance Fields
Traditionally, researchers have been required to know the physical properties of a scene to simulate the rendering process for generating photorealistic images from 3D space. While rendering simulations facilitated the synthesis of controllable high-quality images across the 3D scene, the quality of the synthesized image significantly depends on the physical properties involved in the rendering process. For real-world scenes, estimation of the properties which is referred to as ”inverse rendering” is required, but it is difficult to predict them accurately solely depending on 2D observations like images and videos. Although several approaches have been attempted to overcome the challenges, ”neural rendering” has recently emerged as a superior approach integrating deep learning methodologies and graphics rendering approaches, leveraging the outstanding representation capability of deep neural networks.
According to a comprehensive survey [8], which well summarizes the history of early neural rendering, this research area has been regarded as the intersection of generative adversarial networks (GANs) [9] and graphical controllable image synthesis. With the adoption of GANs, neural rendering has been considered as an image-to-image translation problem utilizing given scene parameters and several 3D scene representations, leveraging the insights of conditional GANs similar to Pix2Pix [10]. For example, [11, 12, 13] generate high-quality images with particular scene conditions by transferring scene parameters to the deep neural network. In addition, other works incorporate the intuition of classical graphics modules into GANs to synthesize and control the image outputs utilizing non-differentiable or differentiable modules such as usage of rendered images with dense input conditioning [14, 11, 15], computer graphics renderer [16, 17], and illumination model [18].
Although these researches present realistic neural rendering techniques over the past few years, there has been great transition in paradigm in neural rendering after emergence of the neural radiance fields (NeRF) [19], which directly map 3D spatial location and viewing direction to irradiance solely relying on multi-view images through multi-layer perceptron (MLP) and classical volume rendering method [20]. NeRF implicitly represents the 3D scene with the classical ray tracing methods and shows photorealistic novel view synthesis, but there is still room for improvement in various aspects. NeRF has widely spread to other computer vision and graphics tasks thanks to its simple and intuitive architecture, which attracts huge attention and expands the research fields of neural rendering. To enhance the performance of neural representation itself, several works have represented 3D scenes using another representation to improve training or rendering speed, such as voxel-grid [21], plenoctree [22], decomposed tensorial fields [23], hashgrid [24], plenoxels [25], light fields [26], and 3D gaussians [27]. In addition, its implicit representation capability leads to explosive development of other graphical tasks such as modeling dynamic scenes [28, 29, 30, 31], relighting [32, 33], 3D reconstructions [34, 35], and human avatar [36].
II-B Radiance Fields in Practical Scenarios
There have been a lot of works to apply the neural representation in more pragmatic scenarios as the importance of VR and AR technologies increased, such as fast rendering, efficient sampling on rays, scene editing, denoising, and training from sparse view. Fast rendering, efficient sampling on rays, and scene editing aim to increase the inference speed, enable surface sampling, and deform the trained mesh through various approaches, such as baking [37], depth-guided sampling [38], and surface deformation [39], respectively.
Another dominant area is constructing the NeRF from sparse view images, which is a practical environment considering real-world scenarios. Sparse view images incur the inherent drawback of neural networks in that the network is more likely to be overfitted to the given data distribution. This leads to inconsistent scene geometry in the mapped representations, typically manifested as incorrectly predicted structural information, such as elongated density artifacts in the rendered color and depth images from novel views. Several approaches have mitigated this issue involving additional prior knowledge or out-of-distribution data. InfoNeRF [40] adopts entropy minimization to probability density function (PDF) of density value along the ray density to make the shape of the PDF sharper. RegNeRF [6] utilizes the statistical depth smoothness of real-world geometry [7] on unobserved ray patches to reduce the artifact. Recently, FlipNeRF [5] considers flipped rays on the surface as supplement unseen rays to regularize the scene geometry. In other approaches, some works, such as PixelNeRF [41], and DietNeRF [42], exploit the semantic information extracted from deep image feature extractors to utilize the representative power of neural networks in feature level. FreeNeRF [4] tries to alleviate the overfitting problem based on an optimization perspective, imposing some restrictions on the frequency level.
In addition to sparse view settings, establishing NeRF from degraded images is recently emerging since the ideal training condition in images for NeRF often breaks in real-world scenarios. RawNeRF [43] denoises the internal noise of the camera sensor to construct high dynamic range (HDR) radiance fields from dark raw images and controls the camera exposure. Similarly, NaN [44] deals with burst noise in images, generating denoised images based on IBRNet [45], which is another image-based rendering approach. For more practical applications of NeRF in real-world, DeblurNeRF [1] firstly attempts to deal with two types of blur degradation in images, blur from camera motion and defocus, constructing deblurred neural radiance fields (DeRF) from only blurry images. They imitate the blurring process integrating the concept of blind deblurring in image deblurring with the NeRF system, modeling the blur kernel as pixel-wise independent ray transformation and composition weights to approximate the blurring process. Another representative approach is proposed by DP-NeRF[2], which imposes physical consistency across the images by modeling the blur kernel as the 3D rigid transformation of rays depending on each view, to approximate the actual blurring process in the camera more precisely. Recently several approaches [46, 47, 3] are also proposed in succession, attempting to improve the quality of the constructed DeRF. One of the most actively researched areas among those mentioned earlier is NeRF from blurry images, which often occurs when users take pictures with their own devices.
However, as we mentioned in Section I, the experimental setup of using only 20$\sim$30 blurry images, as in previous studies, is not practical. If we assume a scenario where users only have access to blurry images, it is more realistic to consider that only a few images are available for a specific scene and all of those images are blurry. Therefore, we propose a more practical scenario by combining DeRF and the sparse view setting, thereby enhancing real-world applicability.
III preliminary
III-A Deblurred Neural Radiance Fields
Neural radiance fields (NeRF) is parameterized MLPs for mapping continuous 3D location $\textbf{x}=(x,y,z)$ to volumetric density $\sigma$ and view-dependent radiance color $\textbf{c}=(r,g,b)$. It is formulated as an approximated universal function $\textbf{F}_{\Theta}:(\gamma_{\textbf{x}}(\textbf{x}),\gamma_{\textbf{d}}(\textbf{d}))\rightarrow(\textbf{c},\sigma)$, where $\Theta$ and $\textbf{d}=(\phi,\theta)$ denote the parameters of the NeRF MLPs and viewing direction of ray, respectively. The function $\gamma$ is a positional encoding function that maps each input x and d to a high dimensional encoded feature, which is generally defined as a concatenation of frequency-adjusted sinusoidal function as Eq. 1.
$\gamma(\textbf{x})=[\textbf{x},\sin(\textbf{x}),\cos(\textbf{x}),...,\sin(2^{f}\pi\textbf{x}),\cos(2^{f}\pi\textbf{x})],$ (1)
where $f=\{0,...,m-1\}$ denotes frequency band with maximum frequency value $m$. Hereafter, we abbreviate the encoding function and represent the function of the NeRF as
$F_{\Theta}(\textbf{x},\textbf{d})\rightarrow(\textbf{c},\sigma).$ (2)
NeRF is trained with pixel-wise color supervision from multi-view input images to optimize the MLPs by predicting each pixel color $\hat{C}$ based on volumetric rendering [48] with the samples along the generated ray r from paired camera parameters. For given ray origin o and viewing direction d along a pixel $p$, the samples along the ray r are evenly divided to $N$ intervals to generate coarse samples with stratified sampling. The samples are defined as $\textbf{r}_{i}=\textbf{o}+t_{i}\textbf{d}$ in near-to-far bounded partitions $[t_{n},t_{f}]$ as shown in Eq.3, where $i$ indicates i-th sample and $t$ denotes the distance from ray origin.
$t_{i}\sim\mathcal{U}\left[t_{n}+\frac{i-1}{N}\left(t_{f}-t_{n}\right),t_{n}+\frac{i}{N}\left(t_{f}-t_{n}\right)\right].$ (3)
Following the [48], the coarse pixel color $\hat{C}_{c}$ is rendered from estimated color $\textbf{c}_{i}$ and density $\sigma_{i}$ of each sample $\textbf{r}_{i}$ as
$\hat{C}(\textbf{r})=\sum_{i=1}^{N}w_{i}\textbf{c}_{i}=\sum_{i=1}^{N}T_{i}\left(1-exp(-\sigma_{i}\delta_{i})\right)\textbf{c}_{i},$ (4)
where $T_{i}=exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ and $\delta_{i}=t_{i+1}-t_{i}$ indicate transmittance of each sample along the ray and distance between adjacent samples, respectively. Hierarchical volume sampling is conducted again utilizing normalized weights as probability density function (PDF) from $w_{i}$ as $\hat{w}_{i}=w_{i}/\sum{w_{j}}^{N_{c}}_{j=1}$ and fine rendered pixel color $\hat{C}_{f}$ is produced through above rendering process again. Coarse- and fine-rendered color is supervised from the true pixel colors from input images through L2-norm as
$\mathcal{L}_{recon}=\sum_{\textbf{r}\in\mathcal{R}}\left[\lVert\hat{C}_{c}(\textbf{r})-C(\textbf{r})\rVert^{2}_{2}+\lVert\hat{C}_{f}(\textbf{r})-C(\textbf{r})\rVert^{2}_{2}\right],$ (5)
where $\mathcal{R}$ is the set of rays in each batch and C(r) is ground truth RGB colors for ray r.
However, the above loss can not be applied to train the DeRF, since there is no true pixel color for training the NeRF in the DeRF environments. To solve this problem and construct DeRF, [1, 2] build additional MLPs for predicting the blur kernel in front of the NeRF to imitate the traditional blind blurring process, which is shown as Eq.6.
$\hat{B}_{p}=\hat{C}_{p}*h_{p},$ (6)
where $p$, $\hat{B}$, *, and $h$ indicate the target pixel, expected blurred color, convolution operator, and blur kernel, respectively. Hereafter, we abbreviate the $p$ for clarity. The expected blurred color $\hat{B}$ is composited from $n$ rendered pixel colors $\hat{C}_{q}$ induced from modeled rays that approximate the blurring process as Eq.7.
$\hat{B}=\sum_{q\in\mathcal{B}(p)}m_{q}\hat{C}_{q},\indent where~{}\mathcal{B}(p)=\{1,\dots,n\},$ (7)
where $m$ and $\mathcal{B}$ denote composition weights and the set of indices of the approximated blurring rays with respect to pixel $p$, respectively. Note that, the number of $\mathcal{B}(p)$ is $n$, which is a hyper-parameter that decides the approximation quality of discrete transformation for blur process. Finally, DeRF is trained with the color reconstruction loss on blurred colors as
$\mathcal{L}^{B}_{recon}=\sum_{\textbf{r}\in\mathcal{R}}\left[\lVert\hat{B}_{c}(\textbf{r})-B(\textbf{r})\rVert^{2}_{2}+\lVert\hat{B}_{f}(\textbf{r})-B(\textbf{r})\rVert^{2}_{2}\right],$ (8)
where $\hat{B}_{c}$, $\hat{B}_{f}$, and $B$ are expected coarse, fine, and ground truth blurred color of the ray r, respectively.
The blur kernels are representatively modeled as a different type of transformation in each paper, [1, 2], which we will describe in the next paragraph. After the joint training of the approximated blur kernel and NeRF simultaneously, they can render the clean neural radiance fields, which we refer to as deblurred neural radiance fields (DeRF), by evaluating only NeRF thanks to the theoretical basis of blind deblurring.
Figure 1: Simple illustration for different blur kernel modeling of (a) Deblur-NeRF [1] and (b) DP-NeRF [2]. The main difference between the two kernels is the consistency between transformed rays induced from the blur kernel. Unlike Deblur-NeRF [1], which independently predict blurring rays for each pixel, DP-NeRF [2] predicts camera motion that makes the blur of each image and shares the motion across the entire pixels in the same image.
III-B Blur Kernels in Deblurred Neural Radiance Fields
The core difference of the blur kernels between the two papers is a consideration of 3D consistency across the entire pixels in each image as we describe in Figure 1. Deblur-NeRF [1] designs the blur kernel as a deformable sparse kernel (DSK), which consists of the $n$ number of transformations depending on embedded latent features from each view and location of image pixels. The transformation is formulated as the 3D vector of ray origin and 2D vector of pixel coordinate, which is initialized within $N\times N$ window on the image plane as
$(\Delta\textbf{v}_{o},\Delta\textbf{v}_{T},m)_{q}=G_{\Phi}(\chi,\textbf{l}_{s}),\indent where~{}q\in\{1,...,n\}.$ (9)
For $G_{\Phi}$, MLP with parameter $\Phi$, inputs are $\chi$ and $\textbf{l}_{s}$, which are latent embedded pixel coordinates and scene information, respectively. Here, $\chi$ is defined as randomly initialized canonical coordinates within a specific small range and $s$ indicates the specific scene. The 3D vector $\Delta\textbf{v}_{\textbf{o}}$ and $\Delta\textbf{v}_{T}$ transform given ray $\textbf{r}=\textbf{o}+t\textbf{d}$ to generate transformed rays imitating blurring process as Eq.10.
$\textbf{r}_{q}=(\textbf{o}+\Delta\textbf{v}_{\textbf{o};q})+t\textbf{d}_{q},$ (10)
where $\textbf{d}_{q}$ is transformed ray direction by applying $\Delta\textbf{v}_{T;q}$ to pixel coordinates to move the endpoint of the ray. Then blurred color $\hat{B}$ of the target pixel is composited by weighted summation of each rendered color $\hat{C}_{q}$ and $m_{q}$ as Eq.7.
However, DP-NeRF [2] argues that the pixel-wise independent optimization of the blur kernel in [1] incurs inconsistency in 3D geometry and appearance. They utilize the physical intuition of actual blurred image acquisition in the camera process as an additional prior for the DeRF, to impose the constraints for constructing radiance fields while preserving 3D consistency. To directly model the actual camera motion as a 3D rigid transformation, they introduce scene-wise $SE(3)$ fields inspired by [29], [49] and Rodrigues’ formula [50]. Scene-wise rigid transformation of the camera is formulated by estimated screw axis $S_{s}\in\mathbb{R}^{6}$ through MLPs depending on only scene information as Eq.11.
$(S_{s;q},m_{s})=(r_{s},v_{s},m_{s})_{q}=T_{\Psi}(\textbf{l}_{s}),\indent where~{}q\in\{1,...,n\},$ (11)
where $T_{\Psi}$ and $\textbf{l}_{s}$ denote MLP with parameter $\Psi$ and latent embedded scene information, respectively. The predicted $r_{q}$ and $v_{q}$ of $S_{q}$ are converted to rotation matrix $e^{r_{q}}$ and translation matrix $\textbf{p}_{q}$ by formulas of [50] and [49], respectively. Note that, we abbreviate specific scene indicator $s$ for clarity. Hence, transformed rays are formulated as the rigid transformation of the rays as Eq.12.
$\textbf{r}_{q}=e^{S_{q}}\textbf{r}=e^{r_{q}}\textbf{r}+\textbf{p}_{q}.$ (12)
The blurred color $\hat{B}$ of the target pixel is also composited by weighted summation of each rendered color $\hat{C}_{q}$ and $m_{q}$ as same as Eq.7. In addition to modeling the blurring process with rigid blur kernel (RBK), [2] proposes an adaptive weight proposal network (AWP) based on the internal correlation between transformed rays and motion axis to predict the adaptive composition weights $\tilde{m}_{q}$ for each pixel, which complements the effect on the blur derived from the depth difference.
In this paper, we present a novel approach that includes geometric constraints and a perceptual prior through extensive experiments based on these two types of blur kernels from Ddeblur-NeRF [1] and DP-NeRF [2], which can be regarded as flexible and rigid kernels, respectively. The conducted experiments demonstrate the effectiveness of Sparse-DeRF on both kernels. Furthermore, the results also reveal a trade-off for the flexibility of the kernels, depending on the specific properties inherent in the diverse scenes.
IV Sparse-DeRF
We find that it is not effective to construct the radiance fields based on naive NeRF, Deblur-NeRF [1], or DP-NeRF [2] from sparse view setting as we present in Section V-E. Although the DeRF usually recovers the high-frequency detail better than the naive NeRF in given views, it has become more fragmented in novel view synthesis, generating inaccurate scene geometry due to the complex joint optimization. The geometric error is represented as mapped RGB textures resembling painted walls in near or far depth and elongated density artifacts, which reveal challenges associated with accurate depth value prediction. However, existing representative regularization for the NeRF from sparse view [6] can not regularize complex joint optimization of the DeRF. We present experimental results in Section V-F1, which shows the difficulty of the previous regularization technique on the DeRF from sparse view. Hence, here we describe our method for regularizing optimization of the DeRF from sparse view to alleviate spatial ambiguity, which consists of two geometric constraints and a perceptual prior. Figure 2 illustrates the overall architecture of the Sparse-DeRF in detail, where (a),(b), and (c) of the Figure 2 indicate each main component of the Sparse-DeRF, respectively. Geometric constraints consist of surface smoothness (SS) and modulated gradient scaling (MGS) as we describe in Section IV-A and Section IV-B. A perceptual prior consists of a perceptual distillation (PD) which is described in Section IV-C.
Figure 2: Overall architecture of the Sparse-DeRF. Each color of integrated ray $\textbf{r}_{io}$ describes where the rays come from. $\hat{d}_{\Theta}(\textbf{r}_{i\sim i+k,j\sim j+k})$ indicates the rendered depth patch with $k\times k$ size. Note that, $k$ is set to $S_{ptc}$ as we described in the paper. The main component of the Sparse-DeRF consists of (a) surface smoothness (SS), (b) perceptual distillation (PD), and (c) modulated gradient scaling (MGS), which are indicated as gray, pink, and green colored components.
Figure 3: Simple illustration of unobserved rays in our method, which consists of unseen rays and hidden rays. Two types of rays are independently defined. Unseen rays remain unchanged during training, but hidden rays change during the training since it is derived from the learned blur kernel. Note that, hidden rays can be derived from both types of blur kernel we utilized.
IV-A Surface Smoothness on Integrated Unobserved Rays
Inspired by the statistical tendencies of real-world geometry, piece-wise smoothness is adopted as a depth smoothness regularization on small rendered patches in RegNeRF [6]. This regularization can also be interpreted as imposing surface smoothness constraint, which is applied to the rendered depth obtained from unseen rays that are defined as rays not observed in the training inputs. The unseen rays are generated from possible camera locations sampled within a limited sample space, constrained by target poses for rendering during test time. Similar to the method introduced in [6], we adopted an approach to alleviate spatial ambiguity by utilizing information from unobserved rays. However, we propose additionally leveraging new unobserved ray information that can only be derived from the blur kernel to stabilize the simultaneous optimization of blur kernel and radiance fields. Firstly, we utilize unseen rays as one of the integrated unobserved rays to ameliorate spatial ambiguity following the [6]. For known set of camera poses $\{\textbf{P}^{i}_{target}\}_{i}$, where $\textbf{P}^{i}_{target}=[\textbf{R}^{i}_{target}|\textbf{t}^{i}_{target}]\in SE(3)$, sampled camera pose $S_{\textbf{P}}$ for unseen rays is formulated from camera location $S_{\textbf{t}}$ and rotation $S_{\textbf{R}}$ in limited sample space as
$\begin{split}S_{\textbf{t}}&=\{\textbf{t}\in\mathbb{R}^{3}|\textbf{t}_{min}\leq\textbf{t}\leq\textbf{t}_{max}\},\\ S_{\textbf{R}}|\textbf{t}&=\{\textbf{R}(\bar{\textbf{p}}_{u},\bar{\textbf{p}}_{f}+\epsilon,\textbf{t})|\epsilon\sim\mathcal{N}(0,0.125)\},\end{split}$ (13)
where $\textbf{t}_{min}$, $\textbf{t}_{max}$, $\bar{\textbf{p}}_{u}$, and $\bar{\textbf{p}}_{f}$ indicate $min(\{\textbf{t}^{i}_{target}\})$, $max(\{\textbf{t}^{i}_{target}\})$, the normalized mean over the up axes of all target poses, and the mean focus point by solving a least squares problem, respectively. $\textbf{R}(\cdot,\cdot,\cdot)$ and $\epsilon$ indicate camera rotation matrix to make the sampled camera roughly focus on a central point of a scene and a small jitter value added to the focus point, respectively. To the end, sampled camera poses $S_{\textbf{p}}$ is formulated as
$S_{\textbf{P}}=\{[\textbf{R}|\textbf{t}]|~{}\textbf{R}\sim S_{\textbf{R}}|\textbf{t},\textbf{t}\sim S_{\textbf{t}}\}.$ (14)
In addition to leveraging previously unseen rays, we employ hidden rays derived from the characteristics of the estimated blur kernel, harnessing supplementary information exclusively presented in blurry inputs. The blur kernel, denoted as $h$ in Eq. 6, generates the ray transformation to approximate the color composition process of the blur, regardless of the kernel types, as demonstrated in Deblur-NeRF [1] and DP-NeRF [2]. Motivated by the commonality in the color composition process across various blur kernels, the kernel-induced transformed rays $\textbf{r}_{q}$ are introduced as additional unobserved rays, which are referred to as hidden rays, to enforce depth smoothness constraint. As the hidden rays $\textbf{r}_{q}$ are not directly presented in the training data for the DeRF system, they serve as supplementary out-of-distribution data for imposing depth smoothness, similar to the aforementioned unseen rays. In addition, the incorporation of depth smoothness on hidden rays effectively addresses geometric inconsistency across the 3D space of blurry images within the specified training view. This capability stems from the broader coverage of the estimated hidden rays, facilitated by the implicit inclusion of camera motion information in blurry images. Therefore, our integrated unobserved rays for depth smoothness are defined as an integrated set of unseen rays and hidden rays as
$\textbf{r}_{iu}=\{\textbf{r}_{q},\textbf{r}_{us}\},~{}\indent where~{}\textbf{r}_{q}\sim h(\textbf{r}),~{}\textbf{r}_{us}\sim S_{\textbf{p}},$ (15)
where $r_{iu}$ and $r_{us}$ denote integrated unobserved rays and unseen rays, respectively. For more intuitive understanding, the $\textbf{r}_{q}$ and $\textbf{r}_{us}$ are illustrated in Figure 3.
For applying depth smoothness constraint on $r_{iu}$, the expected depth of $r_{iu}$ is computed following Eq.16 as same as previous NeRF works.
[…]
Then the depth smoothness loss is reformulated by adding color-dependent weighted depth smoothness from [6] as
$\mathcal{L}_{ss}\{\textbf{r}_{iu}\}$ $=\sum_{\textbf{r}\in\textbf{r}_{iu}}\sum^{S_{ptc}-1}_{i,j=1}\Bigl{[}\Delta\hat{d}_{x}(\textbf{r}_{i,j})+\Delta\hat{d}_{y}(\textbf{r}_{i,j})\Bigl{]},$ (17)
where $\Delta\hat{d}_{x}(\textbf{r}_{i,j})$ and $\Delta\hat{d}_{y}(\textbf{r}_{i,j})$ indicates horizontal and vertial weighted depth difference as Eq.18, respectively.
$\begin{split}\Delta\hat{d}_{x}(\textbf{r}_{i,j})&=\omega_{i+1,j}(\hat{d}(\textbf{r}_{i,j})-\hat{d}(\textbf{r}_{i+1,j}))^{2}\\ \Delta\hat{d}_{y}(\textbf{r}_{i,j})&=\omega_{i,j+1}(\hat{d}(\textbf{r}_{i,j})-\hat{d}(\textbf{r}_{i,j+1}))^{2},\end{split}$ (18)
where $\omega_{i+k,j+1}=exp\big{(}-\big{(}\hat{C}(\textbf{r}_{i,j})-\hat{C}(\textbf{r}_{i+k,j+l})\big{)}^{2}\big{)}$ indicates pixel-wise color difference weight.
Figure 4: The comparison between ours modulated gradient scaling function with $\hat{J}(\delta_{\textbf{s}_{i}})$ and the previous function $J(\delta_{\textbf{s}_{i}})$ proposed by [51], with respect to the ray distance $\delta_{\textbf{s}_{i}}$. In the table, the values of $\delta_{\textbf{s}_{i}}$ and $min(1,J(\delta_{\textbf{s}_{i}}))$ on the x-axis and y-axis represent the ray distance from the camera origin and gradient scaling value, respectively. The $J(\delta_{\textbf{s}_{i}})$ of [51] is represented as $x^{2}$ with the black colored line for clarity. The graphs illustrate that our modulated function $\hat{J}(\delta_{\textbf{s}_{i}})$ can cover the diversity in the arrangement of scene components, exhibiting various shapes of the function depending on magnitude $\rho$ and period $\eta$.
IV-B Modulated Gradient Scaling
Although previous surface smoothness alleviates the spatial ambiguity in the 3D scene, it is still hard to grasp accurate geometry due to the inherent drawback of the NeRF sampling strategy and casted volume occupancy as [51] argued. Following [51], the optimization of the NeRF often fails, generating the density artifact in the near-depth region due to the disproportionate gradient backpropagation induced from the imbalanced volumetric occupancy of the samples on the ray. [51] alleviates the limitation introducing gradient scaling that reduces the propagated gradient $\nabla\textbf{s}_{i}$ of each $i$-th sample $\textbf{s}_{i}=(s_{x},s_{y},s_{z})$ on ray r according to the distance $\delta_{\textbf{s}_{i}}$ from ray origin o as
$\nabla\hat{\textbf{s}}_{i}=\nabla\textbf{s}_{i}\times min(1,J(\delta_{\textbf{s}_{i}})),~{}where~{}\delta_{\textbf{s}_{i}}=|\textbf{s}_{i}-\textbf{o}|,$ (19)
where $\nabla\hat{\textbf{s}_{i}}$ and $J$ indicate scaled gradient value for sample $s_{i}$ and scaling function. The scaling function $J$ is strictly formulated as a squared function as
$J(\delta_{\textbf{s}_{i}})=(\delta_{\textbf{s}_{i}})^{2}.$ (20)
Note that, $\nabla\textbf{s}_{i}$ indicates the gradients of per-point characteristics such as RGB color c and density $\sigma$.
However, we experimentally found that the limitation is more prominent in sparse view settings since there is less available diversity of viewing direction, which means the projective geometry and epipolar geometry do not properly work for NeRF optimization. Therefore, the gradient scaling seems to be more necessary to the radiance fields from the sparse view setting. Although we tried to apply the technique to the Sparse-DeRF, it does not properly works as demonstrated in the appendix. The reason is that the fixed square function of $J$, which is lower bounded as $1$, does not cover the non-linear parametrized space such as normalized device coordinates (NDC), which is commonly used as well as our work. Another reason is that a strictly fixed shape of the function cannot cover the arrangements of the scene components, which are usually different across the scene even in the same dataset. [51] briefly mentioned the determinant of the jacobian as an additional scaling factor for the value of $J$, but they did not experiment on it. Moreover, even if the additional scaling factor were applied, the shape of the scaling function would remain unchanged and simply in the form of the square function, which does not allow for flexible adaptation to the arrangement of scene components. Therefore, we modulate the shape of the scaling function $J$ to adaptively reflect the scene arrangements and be suitable for NDC.
Our novel gradient scaling function $\hat{J}$ is designed based on three conditions. First, the function should increase from zero at the camera origin, which is a critical condition to avoid the local minima in the initial training phase we mentioned before. Second, the function should be not zero in the far distance, which is set to $1$ in our NDC environment, to ensure the NeRF training. Finally, the function is designed to increase and decrease only once within a given depth range, which makes the function not fluctuate, since it is an intuitively reasonable scenario considering the goal of MGS that alleviates incorrect density mapping in near distance. In addition to the above conditions for proper shape of the scaling function, we further consider designing the function shape when the location of the main objects is focused on the center of the scene and density mapping error that is represented as a painted wall of near or far depth. Following the conditions, the proposed modulated gradient scaling (MGS) function $\hat{J}$ is formulated as
$\hat{J}(\delta_{\textbf{s}_{i}})=\rho\Big{(}\sin\big{(}\eta\pi(\delta_{\textbf{s}_{i}}+\frac{3}{2\eta})\big{)}+1\Big{)},$ (21)
where $(1\leq\rho\leq 10)$ and $(\frac{1}{2}\leq\eta<2)$ denote magnitude and period of sinusoidal function, respectively. However, in contrast to [51], the distance range of $\delta_{\textbf{s}_{i}}$ is restricted to $\delta_{\textbf{s}_{i}}\in[0,1]$ since we use NDC for our dataset.
To apply gradient scaling both in the near and far regions while minimizing the scaling effect in the center of the scene, we adopt the sinusoidal function shape as the foundation for our MGS. Furthermore, the second condition determines the maximum value of $\eta$ as 2 to ensure the scaling value of the proposed function in the far region does not fall to or below 0. The function shape in the left top image of Figure 4 shows the characteristics of the proposed scaling function that we described above. We shows the difference between the scaling value from $min(1,J(\delta_{\textbf{s}_{i}}))$ and $min(1,\hat{J}(\delta_{\textbf{s}_{i}}))$, according to the various $\rho$ and $\eta$ values in Figure 4.
Figure 5: Illustration of our perceptual distillation. Perceptual distillation transfers the information of pre-deblurred texture by applying the perceptual loss to the pre-deblurred color patch $\bar{C}_{ptc}$ and rendered color patch $\hat{C}_{ptc}$, which is rendered from patch-wise sampled rays $\textbf{r}^{pd}_{ptc}$ in same pixel location. Note that $\Theta_{D}$ and $\mathcal{E}$ are pre-trained image deblurring network and a shared pre-trained image feature extractor, respectively.
IV-C Perceptual Distillation
In contrast to the previous NeRF-related works in sparse view settings, the Sparse-DeRF environments enable to use the off-the-shelf image processing algorithms, such as deep learning-based image deblurring networks, due to the degradation of the given images. In addition to improvements from a geometric perspective, we aim to enhance the detailed textures of DeRF utilizing the advantages of existing image processing modules, thereby achieving high fidelity.
However, it is not possible to directly utilize the pre-deblurred images as additional pixel-wise color supervision, due to the lack of 3D consistency. This inconsistency occurs due to the ill-posed property of image deblurring and independent deblurring processing across multi-view images, which generates incoherent deblurred results. In the appendix, we additionally address this issue and present the qualitative comparison of pre-deblurred images and reference images, which are estimated from the DP-NeRF [2] trained with the full view, to reveal the geometric inconsistency issue.
To overcome the intrinsic inconsistency of pre-deblurred images, we address the pre-filtered images as a perceptual prior, which transfers the deblurred texture information by extracting features from rendered patches and deblurred images with a pre-trained feature extractor. Figure 5 simply illustrates the pipeline of our perceptual distillation module.
Specifically, we first generate deblurred images $\bar{I}$ for blurry training images $I$ by exploiting a pre-trained image deblurring network $\Theta_{D}$ to prepare the feature extraction as
$\bar{I}=\Theta_{D}(I)$ (22)
Next, we additionally sample the $S^{pd}_{ptc}\times S^{pd}_{ptc}$ size of patch rays $\textbf{r}^{pd}_{ptc}$ and corresponding deblurred image patch $\bar{C}_{ptc}$ from $\bar{I}$ on training views. Then we extract the abundant deblurred features from the rendered color patch $\hat{C}_{ptc}$, which is rendered from $\textbf{r}^{pd}_{ptc}$, and pre-deblurred images $\bar{I}$ using a shared pre-trained image feature extractor $\mathcal{E}$ as
$\hat{\varphi}_{ptc}=\mathcal{E}(\hat{C}_{ptc}),\indent\bar{\varphi}_{ptc}=\mathcal{E}(\bar{C}_{ptc}),$ (23)
where $\hat{\varphi}_{ptc}$ and $\bar{\varphi}_{ptc}$ indicate extracted features from each color patches, respectively. Note that, the color patch $\hat{C}_{ptc}$ is rendered by forwarding the NeRF MLPs without blur kernel to perceptually transfer the pre-filtered texture information to the implicit clean radiance fields, which is indicated as patch radiance in Figure 2. Then we apply the perceptual loss $\mathcal{L}_{pd}$ [52] to distill the feature information as
$\mathcal{L}_{pd}=||\hat{\varphi}_{ptc}-\bar{\varphi}_{ptc}||^{2}_{2}~{}.$ (24)
Our final loss function is a weighted composition of the proposed losses as
$\mathcal{L}_{final}=\mathcal{L}^{B}_{recon}+\lambda_{ss}\mathcal{L}_{ss}+\lambda_{pd}\mathcal{L}_{pd},$ (25)
where $\lambda_{ss}$ and $\lambda_{pd}$ denote weights for each loss, which are equally set to $0.01$ in our experiments.
V Experiments
V-A Dataset
Sparse-DeRF has experimented with a forward-facing scene dataset proposed by Deblur-NeRF [1], which includes 5 synthetic and 10 real scenes. In particular, we use only the camera motion blur dataset since our goal is to alleviate the blur from camera motion in the DeRF from sparse view settings. The dataset consists of multiple view images and paired camera poses calibrated by using COLMAP [53, 54]. For the sparse view setting, we manually select the 2, 4, and 6 images as training datasets for all scenes, so that the entire space covered by each view is as wide as possible. In addition, to ensure reasonable learning of radiance fields, we select views with visible spaces that overlap as little as possible, while avoiding excessively extreme blur magnitudes. Note that, it is an inherent property of the dataset that blur magnitudes of selected views can be different according to each scene, which means that each scene has a different level of learning difficulty. We attach the selected image indices of all scenes used in our experiments in the appendix for fair comparison in future research.
V-B Experimental Sparse View Setting
Scenarios for the Sparse-DeRF are assumed to be three kinds of settings, which are the DeRF from 2-view, 4-view, and 6-view settings unlike existing common sparse-view NeRFs, which usually use 3-view, 6-view, and 9-view settings. The reason is that when using more than 8 images, the joint optimization problem occurs more frequently under 9-view settings as shown in Figure 6. The figure presents the graph of experiments in the Decoration scene where we varied the number of input sparse views from 2-view to 10-view. We also attach the quantitative results in the appendix due to the page limitation. Experimental results reveal that our sparse-view experimental settings are valid since the RegNeRF [6] with DP-kernel shows poor performances under the 9 views.
V-C Implementation Details
Spare-DeRF is implemented and modified based on the published official code of DP-NeRF [2] using the two types of blur kernels of DP-NeRF [2] and Deblur-NeRF [1]. For a fair comparison, the number of blurring rays is set to $5$ for the default setting as same as previous works [1, 2]. We set the other settings for the blur kernels following the default parameters of each work. For the NeRF optimization, we use $64$ coarse and $64$ fine samples per ray with a batch size of 1024 rays, exploiting Adam [55] optimizer with default parameters. In addition, exponential weight decay is applied from $5\times 10^{-4}$ to $8\times 10^{-5}$ for learning rate scheduling. We train the DeRF for $20k$, $40k$, and $60k$ iterations in 2-view, 4-view, and 6-view settings, respectively. For patch-wise sampling, we set the size of the patch $S_{ptc}$ and $S^{pd}_{ptc}$ as $8$ and $64$, respectively. However, the generating method of unseen rays in surface smoothness constraint is especially considered for fair comparison across the extensive experiments. In particular, we generate the unseen rays from fixed views, which are evenly selected from the rest of the training rays, to remove the randomness of the unseen ray generation for experimental analysis and fair comparison. Note that, the rest of the training rays are not included in the training or test view. We demonstrate that it is a reasonable choice since the selected views are still in the sample space we defined in Section IV-A. In addition, we attach the visualization of this issue in the appendix to show the rationality. The hyper-parameters of the modulated gradient scaling function for each scene, magnitude $\rho$ and period $\eta$, are attached in the appendix in detail. For perceptual distillation, MPRNet [56] is utilized as a pre-trained image deblurring network. We select the VGG19 [57] as a pre-trained image feature extractor $\mathcal{E}$ since it is widely exploited as an image feature extractor in image-based computer vision.
V-D Evaluation Metrics
Our experimental results for the synthetic and real datasets are evaluated in three quantitative metrics and qualitative comparisons between rendered images through a novel view synthesis task. Consistent with prior research, we employ widely utilized evaluation metrics to compare the synthesized images with corresponding ground truth images: the peak signal-to-noise ratio (PSNR), the structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS) [58]. These metrics assess the relative sharpness, structural similarity, and perceptual quality of the generated images, respectively. In addition, we encourage readers to refer to the supplementary video for a more comprehensive and detailed presentation of the results.
Figure 6: Graph of the experimental results of Decoration scene from 2-view to 10-view settings. Figure (a), (b), and (c) indicate PSNR, SSIM, and LPIPS, respectively. The graph reveals that the network suffers joint optimization problems under the 9-view setting, which makes our experimental setting plausible.
V-E Evaluation
TABLE I: Average results of novel view synthesis for both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings. Each color shading represents the best, second best, and third best result for each experimental setting, respectively.
[ 2-view ] [ 4-view ] [ 6-view ]
[ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ]
 PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 15.11 .2999 .5578 14.38 .2635 .6004 20.02 .5327 .4000 18.98 .4860 .4481 21.65 .5985 .3638 20.63 .5513 .4014
MPR [56] + NeRF 15.16 .3006 .5595 14.38 .2594 .6019 20.00 .5381 .3956 18.89 .4829 .4484 21.72 .5999 .3629 20.60 .5513 .4010
Deblur-NeRF [1] 15.14 .2884 .5330 14.41 .2506 .5921 19.99 .5199 .3499 18.89 .4761 .4151 23.12 .6798 .2386 21.36 .6003 .3163
DP-NeRF [2] 15.06 .2827 .5389 14.36 .2506 .5904 19.84 .5336 .3075 18.77 .4582 .4175 23.68 .7036 .1998 21.68 .6137 .2992
RegNeRF [6] (No kernel) 14.60 .2849 .5869 15.49 .2997 .5888 18.39 .4600 .4704 18.44 .4326 .4852 19.69 .5249 .4165 19.65 .4790 .4583
RegNeRF [6] (w/DP-kernel) 13.24 .2162 .6062 12.76 .1836 .6447 16.44 .3657 .4826 13.59 .2221 .5887 21.60 .6162 .3046 18.25 .4439 .4326
Sparse-DeRF (w/DN-kernel) - Ours 15.52 .2966 .5291 15.53 .3112 .5515 20.57 .5565 .3354 19.98 .5231 .3871 23.32 .6903 .2379 22.15 .6248 .3030
Sparse-DeRF(w/DP-kernel) - Ours 15.35 .2904 .5242 15.57 .3114 .5467 21.05 .5776 .2975 20.05 .5178 .3736 24.27 .7255 .2044 22.32 .6283 .2907
Figure 7: Qualitative results on ParterreCoffee, and Decoration scene from (a) 2-view, (b) 4-view, and (3) 6-view respectively. (i), (ii), and (iii) denote rendered color and depth images from DP-NeRF [2], RegNeRF [6] with DP-kernel, and Ours(Sparse-DeRF) with DP-kernel.
V-E1 Quantitative Evaluation
We present the quantitative results of Sparse-DeRF for two different types of blur kernels from Deblur-NeRF [1] and DP-NeRF [2], comparing these results with established baseline methods. The effectiveness of our approach is demonstrated in TABLE I, achieving outstanding performance across entire sparse view settings, regardless of the blur kernel employed. In the Sparse-DeRF results, DN-kernel and DP-kernel denote the blur kernels proposed by the Deblur-NeRF [1] and DP-NeRF [2], respectively. MPR$+$NeRF denotes the naive NeRF model trained solely on color supervision from deblurred images by MPRNet [56], utilizing the reconstruction loss $\mathcal{L}_{recon}$ of Eq.5. The results of MPR$+$NeRF reveal interesting observations and marginal improvements in radiance fields when only employing pre-deblurred images as direct color supervision for training. This tendency emphasizes the 3D inconsistency across the pre-deblurred images, as discussed in Section IV-C. In addition, the poor results of the RegNeRF [6] with and without a blur kernel demonstrate that existing regularization faces difficulty in alleviating the complex joint optimization involving both the blur kernel and radiance fields. We further present an analysis of this optimization issue in Section V-F1 with detailed experimental results. In contrast, Sparse-DeRF demonstrates significant enhancements across all evaluation metrics for both types of blur kernels, indicating its superior ability to represent the DeRF with improved visual quality. In particular, our results exhibit more prominent improvements in real-scene scenarios, although there is non-ideal blur degradation, which occurs due to various real environmental factors. For a comprehensive understanding, we provide an extensive ablation study in Section V-F, incorporating results with both types of blur kernels. Additionally, detailed results for all scenes are appended in the appendix, including synthetic and real scenes.
V-E2 Qualitative Evaluation
In Figure 7, we present representative qualitative results on three scenes (Parterre, Coffee, and Decoration) from 2-view, 4-view, and 6-view settings, respectively. The figure depicts the results of novel view synthesis, presenting rendered color and depth images. The figures demonstrate that our model significantly enhances the visual quality of radiance fields in terms of geometric and perceptual fidelity. In addition to the above quantitative results, qualitative results also demonstrate the inconsistency issue of pre-deblurred images and complex joint optimization of the DeRF from sparse view. All the results (ii) of Figure 7 demonstrate that existing representative regularization technique, RegNeRF [6], can not effectively alleviate the optimization issue of the DeRF from sparse view. Furthermore, we encourage readers to view the supplementary videos that emphasize 3D consistency through rendered videos from a spiral camera path.
V-F Ablations
V-F1 Problem Analysis and Motivation
we present the experimental results to describe that it is difficult to jointly optimize the DeRF and naively apply the previous regularization technique of the NeRF to the DeRF from sparse view setting, utilizing the representative existing regularization method, RegNeRF [6]. We attach the experimental results of the RegNeRF [6] from 2-view, 4-view, and 6-view settings with and without the blur kernel to demonstrate the difficulty of the complex joint optimization problem as we mentioned. TABLE II and Figure 8 present quantitative and qualitative results from 2-view, 4-view, and 6-view, respectively. The blur kernel employed is the rigid blur kernel of DP-NeRF [2], which is described as DP-kernel in the table and figures. To help the reader compare the plausible appearance and dense geometry with the proposed Sparse-DeRF, we attach the rendered color and depth images of our model. The results of the figures demonstrate that the integration of the blur kernel involves a straightforward difficulty in optimizing the high-frequency details and the overall scene geometry simultaneously, as indicated by the visual quality of the rendered color and depth images. Although the presence of a blur kernel enables radiance fields to capture high-frequency details, it causes a geometric distortion with the overall wrong density mapping that resembles the fragmented structure of objects or appearance like the painted wall of near- or far-depth regions. These results and analysis demonstrate that naively applying the regularization of the NeRF from sparse view to the DeRF is not effective. In addition, the images (i) and (ii) in Figure 8 demonstrate that although the overall performance is better without the blur kernels, the model still faces difficulty in modeling high-frequency details. This difficulty makes the necessity of the blur kernel optimization still important, leading to a blurry visual quality across the entire scene. Therefore, motivated by the experimental analysis, we propose Sparse-DeRF, the novel regularization method for optimization of the blur kernel and radiance fields simultaneously. The proposed Sparse-DeRF presents high-quality rendered images with dense geometry and detailed high-frequency texture as shown in Figure 8.
TABLE II: Average results for the complex optimization issue in both synthetic and real scenes from 2-view, 4-view, and 6-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
[ 2-view ] [ 4-view ] [ 6-view ]
[ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ]
Blur Kernel  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 14.60 .2849 .5869 15.49 .2997 .5888 18.39 .4600 .4704 18.44 .4326 .4852 19.69 .5249 .4165 19.65 .4790 .4583
RegNeRF [6] DP-kernel 13.24 .2162 .6062 12.76 .1836 .6447 16.44 .3657 .4826 13.59 .2221 .5887 21.60 .6162 .3046 18.25 .4439 .4326
Sparse-DeRF (Ours) DP-kernel 15.35 .2904 .5242 15.57 .3114 .5467 21.05 .5776 .2975 20.05 .5178 .3736 24.27 .7255 .2044 22.32 .6283 .2907
Figure 8: Qualitative results on Pool, Tanabata, and Heron scene for complex optimization issue from 2-view, 4-view, and 6-view settings. (i), (ii), and (ii) denote the rendered images from the RegNeRF [6] with no blur kernel, RegNeRF with DP-kernel, and Sparse-DeRF(Ours) with DP-kernel, respectively.
TABLE III: Ablation experimental results from 2-view, 4-view, and 6-view according to the proposed geometric constraints and perceptual prior, denoted in the table as surface smoothness (SS), modulated gradient scaling (MGS), and perceptual distillation (PD), respectively. We present separate results for our model with both types of kernels, DP-kernel [2] and DN-kernel [1]. Each color shading represents the best, second best, and third best result for each experimental setting, respectively.
[ 2-view ] [ 4-view ] [ 6-view ]
SS MGS PD DP-kernel DN-kernel DP-kernel DN-kernel DP-kernel DN-kernel
 PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Synthetic Scene 15.06 .2827 .5389 15.14 .2884 .5330 19.84 .5336 .3075 19.99 .5199 .3499 23.68 .7036 .1998 23.12 .6798 .2386
14.88 .2772 .5388 15.25 .2828 .5418 19.93 .5411 .3107 19.89 .5201 .3558 24.15 .7263 .1974 23.59 .6795 .2429
15.40 .2857 .5289 15.12 .3025 .5243 20.16 .5425 .3024 19.44 .5295 .3562 23.83 .7147 .2072 24.33 .6832 .2406
15.03 .2796 .5391 14.84 .2702 .5454 19.65 .5237 .3311 19.72 .5104 .3630 23.89 .7150 .2052 23.35 .6941 .2341
14.52 .2647 .5481 15.11 .2889 .5389 20.91 .5812 .2948 20.03 .5294 .3450 24.28 .7279 .2021 23.41 .6941 .2324
15.50 .2954 .5215 15.57 .2979 .5319 19.65 .5164 .3334 19.96 .5222 .3547 23.53 .7088 .2153 22.91 .6654 .2560
15.35 .2904 .5242 15.52 .2966 .5291 21.05 .5776 .2975 20.57 .5565 .3354 24.27 .7255 .2044 23.32 .6903 .2379
Real Scene 14.36 .2506 .5904 14.41 .2506 .5921 18.77 .4582 .4175 18.89 .4761 .4151 21.68 .6137 .2992 21.36 .6003 .3163
14.26 .2414 .5933 14.33 .2495 .5937 19.04 .4760 .4008 18.96 .4839 .4125 22.10 .6214 .2951 21.53 .6044 .3178
15.46 .3035 .5465 15.44 .3037 .5558 19.75 .4980 .3737 19.06 .4907 .4105 22.07 .6212 .2889 21.76 .6101 .3097
14.28 .2490 .5896 14.47 .2549 .5939 19.00 .4724 .4080 19.84 .5118 .3937 21.73 .6082 .3082 21.69 .6087 .3142
14.00 .2587 .5830 14.47 .2563 .5881 19.00 .4760 .4037 18.94 .4861 .4145 21.84 .6105 .3084 21.70 .6129 .3144
15.40 .3009 .5527 15.46 .3073 .5519 19.74 .4986 .3774 19.93 .5188 .3873 $2$6257 .2868 21.88 .6154 .3112
15.57 .3114 .5467 15.53 .3112 .5515 20.05 .5178 .3736 19.98 .5231 .3871 22.32 .6283 .2907 22.15 .6248 .3030
Figure 9: Qualitative ablation results of each component on Girl, Stair, and Coffee scenes from 2-view, 4-view, and 6-view settings, respectively. We attach the rendered depth images of the DP-NeRF [2] trained from full view as reference depth images to help the reader compare the results.
V-F2 Ablation of Each Component
We demonstrate the effectiveness of the Sparse-DeRF’s each component through comprehensive ablation studies, presenting both quantitative and qualitative results. In TABLE III and Figure 9, we present independent quantitative and qualitative results from SS, MGS, and PD, where each component is individually applied to observe the influence of each method. The results include several combinations of the components to reveal the complement effect of each proposed method. The experiments are conducted for entire 2-view, 4-view, and 6-view settings, providing a different performance depending on two types of blur kernels, DN-kernel [1] and DP-kernel [2]. Our model shows superior results in predicting both 3D geometric and appearance details precisely, demonstrating the enhanced evaluation results.
V-F3 Ablation Analysis
Quantitative ablation results demonstrate that our model with full components shows the best results in the real-scene dataset. The results reveal that MGS plays an important role across the entire 2-view, 4-view, and 6-view settings among geometric constraints although each constraint enhances the geometric accuracy. In Figure 9, the importance of MGS is more clearly prominent with qualitative results as color and depth images. If MGS is not applied, we can see that the geometry of the scene is not accurately captured or there are many density artifacts. On the other hand, PD seems to be not effective without geometric constraints in 2-view and 4-view settings as we can figure out in TABLE III and Figure 9. The reason is that the NeRF has more difficulty in predicting the correct geometry with only pre-deblurred images due to the inherent 3D inconsistency, which makes perceptual prior not effective. These difficulties become more severe as the number of views decreases. We can demonstrate that a certain level of accurate geometry should be achieved before applying perceptual distillation.
TABLE IV: Average quantitative comparison results between naive gradient scaling of [51] and our MGS in both synthetic and real scenes from 2-view, 4-view, and 6-view settings. The two types of kernels we utilize in this paper are indicated as DP-kernel and DN-kernel, the kernels of DP-NeRF [2] and Deblur-NeRF [1], respectively. Color shading represents the better result.
[ 2-view ] [ 4-view ] [ 6-view ]
[ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ]
Blur Kernel Gradient Scaling  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel + Naive [51] 14.70 .2483 .5643 14.85 .2860 .5653 18.69 .4321 .4242 19.28 .4909 .4119 22.00 .6002 .3117 21.20 .5877 .3274
DN-kernel + MGS 15.12 .3025 .5243 15.44 .3037 .5558 19.44 .5295 .3562 19.84 .5118 .3937 22.40 .6832 .2406 21.76 .6101 .3097
DP-kernel + Naive [51] 14.01 .2377 .5743 14.95 .2871 .5591 17.25 .3933 .4221 18.85 .4670 .4014 21.55 .6097 .2838 21.49 .6010 .3111
DP-kernel + MGS 15.40 .2857 .5289 15.46 .3035 .5465 20.16 .5425 .3024 19.75 .5980 .3737 23.83 .7147 .2072 22.07 .6212 .2889
Figure 10: Qualitative results on Puppet, Trolley, and Factory scene for comparison between naive gradient scaling [51] and our proposed MGS from 2-view, 4-view, and 6-view settings. Note that we attach the rendered color and depth images from DP-NeRF [2] trained with full view as reference images.
V-F4 Comparison to Naive Gradient Scaling
We present experimental results of quantitative and qualitative comparison between our proposed MGS and naive gradient scaling of [51] from 2-view, 4-view, and 6-view settings in TABLE IV and Figure 10. We compare the effectiveness of our MGS for two types of kernels we utilize in our paper, which are the kernels of Deblur-NeRF [1] and DP-NeRF [2]. The results describe our MGS outperforms the naive gradient scaling in terms of both quantitative and qualitative performance across the entire experimental setting. Specifically, depth images demonstrate that our MGS more effectively helps the model to predict the accurate geometry than naive gradient scaling. Quantitative results for the entire scene are attached in the appendix.
V-F5 Inconsistency of Pre-deblurred Images
As mentioned in Section I and IV-C, image deblurring is conducted independently for each image and presents inconsistent geometry in various regions due to its ill-posed property. In Figure 11, we present the qualitative comparison to reveal the inconsistency of pre-deblurred images across the multi-view training images. Pre-deblurred images are acquired by applying the MPRNet [56]. In addition, we attach the reference images, which are rendered from the DP-NeRF [2] trained with full view, to help readers better understand the inconsistency issue and compare the pre-deblurred geometry to approximated ground truth geometry. Comparing emphasized regions from each image, the pre-deblurred image shows relatively well-restored textures in each image, but the geometry is inconsistently restored and distorted across the multiple views. Such inconsistency adversely affects the learning of radiance fields since it is trained by pixel-wise color reconstruction loss. Furthermore, the performance of image deblurring varies even within a single image depending on the structural complexity of the local region, making it challenging to learn blur kernels. Due to these reasons, directly utilizing pre-filtered images for training radiance fields is difficult. Therefore, perceptual distillation is introduced to transfer only the perceptual texture of pre-deblurred images to the radiance fields.
Figure 11: Qualitative comparison on Girl scene that demonstrates the geometric inconsistency of pre-deblurred images. Figure (a)$\sim$(c) presents input blurry images $I$ for training, pre-deblurred images $\bar{I}$ by MPRNet [56], and rendered images from DP-NeRF [2] trained with full view, respectively. Figure (d)$\sim$(f) shows the emphasized regions of Figure (a)$\sim$(c), which demonstrates the inconsistency issue in detail.
VI Limitations and Discussions
Despite the remarkable enhancement in terms of 3D geometry and appearance, there are still several limitations. The first one is derived from the blur kernel itself, especially the relationship between the type of the blur kernel and the properties of each scene. For example, the performance is more improved with the rigid blur kernel of DP-NeRF [2] in some scenes, but in other scenes, the improvements are greater with the flexible blur kernel of Deblur-NeRF [1]. These kernel-dependent performances are different across the scenes. As we figure out, a flexible kernel leads to reduced space ambiguity but high scene distortion. In contrast, the rigid kernel leads to accurate geometry but suffers difficulty in optimizing the scene where the distances of the objects from the camera in the scene are diverse and some objects are located very close to the camera due to the inherent rigidity. We tried to take advantage of both kernels and design the hybrid kernel to maximize the effectiveness of the Sparse-DeRF, but it didn’t work as we imagined. Constructing the hybrid blur kernel that has rigid and flexible properties can be a promising future research direction regardless of sparse view setting in the deblurred neural radiance fields (DeRF). The second one is that we have to set the proper hyper-parameter for MGS to find the most effective function shape although MGS greatly improves the 3D geometry in the DeRF from sparse view. However, it is difficult to find an ideal function shape according to the arrangement of the object in the scene, especially as we mentioned above. We handle these cases by setting the magnitude $\rho$ as a high value to only ignore the gradient in a very near distance region, which is attached to the appendix as detailed hyper-parameters per each scene. In this sense, this manual setting of hyperparameters is regarded as one of our limitations. We believe the limitation can be alleviated in future research through various methods such as the introduction of learnable parameters for gradient scaling function. Finally, although sparse view setting of blurry inputs is an extremely practical scenario for blurry inputs, it is too hard to enhance the performance of the DeRF in the 2-view setting due to the lack of the scene information included in the input data. The innate challenge of the 2-view setting is that sparse overlapped 3D space usually leads to inaccurate geometry, which is more likely to be mapped to be painted texture on the wall at the near or far depth regions. There is still room to improve the visual quality of the Sparse-DeRF and solve these problems through state-of-the-art generative methods such as diffusion models, which can be a great future direction for constructing the DeRF from sparse view.
VII Conclusion
In this work, we propose the Sparse-DeRF, a novel regularization method for high-quality deblurred neural radiance fields from sparse view settings, which considers more practical real-world scenarios for radiance fields from only blurry images. We propose two geometric constraints that consist of surface smoothness and modulated gradient scaling, which reflect the real-world statistical geometry and alleviate elongated density artifacts in deblurred neural radiance fields system from sparse view. In addition, we propose a perceptual distillation to utilize the pre-deblurred images as a perceptual prior, which enhances the sharp texture on deblurred neural radiance fields. We demonstrate the effectiveness of the Sparse-DeRF that ameliorates the spatial ambiguity and structural distortion of deblurred neural radiance fields by presenting extensive experimental results in 2-view, 4-view, and 6-view settings. As deblurred neural radiance fields have attracted attention across the research fields related to neural rendering, we believe our work presents a way for future research directions since we address the more practical scenarios for deblurred neural radiance fields from blurry images.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2024-00340745) and an Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government [24ZC1200, Research on hyper-realistic interaction technology for five senses and emotional experience]
Appendix A Sample Space Visualization
In Figure 12, we present the visualization of the sample space of the unseen rays from the RegNeRF [6] and our fixed unseen rays, which is used for fair comparison in the paper. Since they randomly sample the camera poses from the sample space in every training, we sample the $50$ unseen camera poses from the sample space to show the approximate coverage of the sample space. As we can see in the Figure 12, the fixed unseen rays, which are used in our experiments still in the coverage of the sample space of the RegNeRF. In addition, training camera poses and fixed unseen training rays do not significantly overlap with test rays, which also does not break the training and testing rule for novel view synthesis. Hence, it is not a problem to use the fixed unseen rays as alternative unseen rays of the RegNeRF. As we mentioned in Section V-C in the main paper, we utilize the fixed unseen rays for training our model to fairly evaluate the performances across the extensive experiments since the randomness of unseen ray generation in the RegNeRF makes it hard to understand the effectiveness of each component.
TABLE V: Image indices of each scene for training Sparse-DeRF.
Real Scene Ball Basket Buick Coffee Decoration Girl Heron Parterre Puppet Stair
$2$-view 1, 12 12, 33 11, 39 3, 10 1, 19 9, 16 11, 35 8, 26 9, 31 13, 26
$4$-view 1, 12, 18, 22 1, 12, 22, 33 5, 11, 20, 39 3, 10, 15, 26 1, 19, 22, 39 2, 9, 16, 32 4, 11, 18, 35 1, 8, 13, 26 9, 13, 21, 31 4, 13, 16, 26
$6$-view 1, 5, 10, 12, 18, 22 1, 8, 12, 17, 22, 33 5, 11, 17, 20, 34, 39 3, 10, 11, 15, 21, 26 1, 14, 19, 22, 27, 39 2, 9, 16, 24, 32, 37 4, 11, 18, 23, 27, 35 1, 8, 13, 17, 26, 28 7, 9, 13, 21, 23, 31 2, 4, 13, 16, 26, 34
Synthetic Scene Cozyroom Factory Pool Tanabata Trolley
$2$-view 2, 17 3, 19 10, 23 1, 7 13, 23
$4$-view 2, 17, 23, 29 3, 14, 19, 33 5, 10, 15, 23 1, 7, 11, 22 7, 13, 23, 31
$6$-view 2, 14, 17, 21, 23, 29 1, 3, 14, 19, 28, 33 1, 5, 10, 15, 20, 23 1, 7, 11, 18, 22, 27 7, 13, 20, 23, 27, 31
TABLE VI: Hyper parameters of modulated gradient scaling (MGS) for entire scenes of the Deblur-NeRF [1] dataset. $\rho$ and $\eta$ denote the magnitude and period of sine function, respectively.
Synthetic Scene Cozyroom Factory Pool Tanabata Trolley
$\rho$ 10.0 10.0 10.0 1.0 10.0
$\eta$ 1.75 1.75 1.75 1.5 1.75
Real Scene Ball Basket Buick Coffee Decoration Girl Heron Parterre Puppet Stair
$\rho$ 1.0 1.0 10.0 1.0 1.0 1.0 1.0 1.0 10.0 1.0
$\eta$ 1.2 0.67 1.75 0.67 0.5 0.5 0.5 0.5 1.75 0.5
Appendix B Additional Implementation Details
B-A Training Scene Indices
For fair comparison in future research, we present the image indices of each scene for training the Sparse-DeRF in TABLE V. The indices are manually selected from the training images of each scene as we mentioned in the main paper. Note that, the indices of the 2-view and 4-view settings are subsets of the 6-view setting.
B-B Parameters for Entire Scenes
In TABLE VI, we present the hyper-parameters of MGS for entire scenes, which consist of period $\eta$ and magnitude $\rho$ of the sine function. As we indicate in Section IV-B in the main paper, we set the magnitude $\rho$ as high value to only ignore the gradient in very near distance regions for scenes such as Buick, Puppet, Cozyroom, Factory, Pool, and Trolley. Please refer to the figure of the function shape depending on the parameters in the main paper.
Appendix C Additional Quantitative Results
C-A Quantitative Results for Entire Scenes
We present the comprehensive experimental results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE VIIVIII, and IX, respectively. The results of the Sparse-DeRF in the TABLE VIIVIII, and IX are representative results as we presented in Section V-E1 in the main paper. In addition, we present more detailed ablation results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XXI, and XII, respectively. As we mentioned before, we also present the separated ablation results according to the type of the kernel from DP-NeRF [2] and Deblur-NeRF [1].
C-B Complex Optimization Problem for Entire Scenes
We present the comprehensive experimental results of the complex optimization problem for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XIVXV, and XVI, respectively. In addition, we also present the quantitative results of the experiments on the Decoration scene that varies the number of sparse views in TABLE XIII as we mentioned in the main paper.
C-C Comparison to Naive Gradient Scaling for Entire Scenes
We present the quantitative comparison results for entire scenes that compare our modulated gradient scaling (MGS) and naive gradient scaling of [51] from 2-view, 4-view, and 6-view settings in TABLEXVIIXVIII, and XIX, respectively.
TABLE VII: Quantitative results of novel view synthesis for the entire scene of synthetic and real scenes obtained from 2-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 14.24 .2186 .6867 21.07 .6127 .3644 18.59 .3575 .4288 10.84 .1517 .6500 10.82 .1591 .6593 15.11 .2999 .5578
MPR [56]+NeRF 14.43 .2250 .6858 21.08 .6111 .3671 18.72 .3676 .4228 11.03 .1456 .6553 10.56 .1535 .6663 15.16 .3006 .5595
Deblur-NeRF [1] 14.14 .2009 .6647 20.58 .5768 .3443 18.37 .3247 .4061 11.56 .1704 .6075 11.05 .1690 .6425 15.14 .2884 .5330
DP-NeRF [2] 14.14 .2091 .6540 20.71 .5752 .3487 18.38 .3276 .4127 11.21 .1482 .6247 10.88 .1536 .6543 15.06 .2827 .5389
RegNeRF [6] (No kernel) 14.57 .2523 .6680 17.13 .4616 .3808 14.23 .1473 .7298 13.01 .2408 .5956 14.04 .3227 .5602 14.60 .2849 .5869
RegNeRF [6] (w/DP-kernel) 14.20 .2214 .6412 18.88 .5136 .3476 13.03 .1003 .6845 9.21 .0815 .6920 10.86 .1644 .6657 13.24 .2162 .6062
Sparse-DeRF (w/DN-kernel) - Ours 14.27 .2200 .6564 20.57 .5675 .3589 19.81 .3330 .4058 11.67 .1827 .5947 11.28 .1800 .6298 15.52 .2966 .5291
Sparse-DeRF (w/DP-kernel) - Ours 14.10 .2116 .6413 18.97 .5206 .3417 20.32 .3488 .4056 12.00 .1943 .5902 11.36 .1769 .6422 15.35 .2904 .5242
Real Scene Ball Basket Buick Coffee Decoration
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 18.45 .4363 .5693 13.32 .2595 .6131 13.49 .2620 .5680 17.76 .5249 .4807 12.25 .1718 .6738
MPR [56]+NeRF 18.69 .4413 .5659 13.09 .2505 .6194 13.31 .2574 .5706 18.13 .5243 .4869 12.41 .1761 .6735
Deblur-NeRF [1] 18.80 .4325 .5521 12.86 .2414 .5947 13.11 .2446 .5609 18.66 .5272 .4696 12.09 .1508 .6722
DP-NeRF [2] 18.22 .4200 .5611 13.24 .2263 .6062 13.43 .2480 .5571 18.15 .5191 .4876 12.34 .1672 .6675
RegNeRF [6] (No kernel) 20.48 .4951 .5398 15.53 .3388 .5403 17.05 .4241 .4361 23.24 .6950 .3505 10.94 .0816 .7545
RegNeRF [6] (w/DP-kernel) 18.58 .4297 .5565 13.49 .2429 .5985 12.51 .2199 .5579 15.71 .4280 .5370 11.15 .0927 .7520
Sparse-DeRF (w/DN-kernel) - Ours 20.07 .4612 .5290 13.62 .2678 .5835 14.38 .3182 .4913 19.51 .6048 .4146 13.09 .2173 .6256
Sparse-DeRF (w/DP-kernel) - Ours 20.09 .4583 .5244 14.05 .2890 .5555 13.78 .2864 .5036 19.67 .6084 .4096 13.11 .2192 .6192
Real Scene Girl Heron Parterre Puppet Stair Average
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 11.06 .2298 .6304 13.69 .1564 .5683 15.20 .2528 .6462 14.34 .2937 .6156 14.27 .0474 .6385 14.38 .2635 .6004
MPR [56]+NeRF 10.80 .1977 .6424 13.51 .1538 .5652 15.45 .2667 .6422 14.22 .2844 .6135 14.15 .0414 .6392 14.38 .2594 .6019
Deblur-NeRF [1] 10.94 .2031 .6436 13.97 .1459 .5499 15.02 .2320 .6431 14.28 .2764 .6036 14.40 .0525 .6313 14.41 .2506 .5921
DP-NeRF [2] 11.14 .2151 .6485 13.63 .1473 .5376 14.70 .2226 .6348 14.32 .2887 .5930 14.38 .0521 .6104 14.36 .2506 .5904
RegNeRF [6] (No kernel) 8.31 .0602 .7545 11.43 .0951 .6687 15.58 .2709 .6510 16.02 .3548 .5739 16.30 .1816 .6182 15.49 .2997 .5888
RegNeRF [6] (w/DP-kernel) 7.86 .0419 .7737 11.57 .0783 .6246 12.54 .1524 .7091 10.67 .1331 .6604 13.49 .0170 .6770 12.76 .1836 .6447
Sparse-DeRF (w/DN-kernel) - Ours 12.85 .3418 .5585 14.30 .1706 .5197 16.61 .2903 .5912 14.91 .3141 .5886 15.97 .1255 .6125 15.53 .3112 .5515
Sparse-DeRF (w/DP-kernel) - Ours 13.11 .3707 .5486 14.17 .1669 .5285 16.19 .2527 .5908 15.25 .3277 .5739 16.30 .1343 .6133 15.57 .3114 .5467
TABLE VIII: Quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 4-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 16.63 .3494 .5861 22.75 .6935 .2903 25.92 .6419 .2495 16.15 .4408 .4635 18.66 .5380 .4104 20.02 .5327 .4000
MPR [56]+NeRF 16.97 .3570 .5789 22.86 .7014 .2895 25.77 .6436 .2465 16.78 .4827 .4361 17.61 .5056 .4272 20.00 .5381 .3956
Deblur-NeRF [1] 17.26 .3740 .5088 25.51 .7767 .1897 23.38 .5068 .2649 15.91 .4209 .4198 17.91 .5213 .3661 19.99 .5199 .3499
DP-NeRF [2] 19.20 .5175 .3813 21.50 .6242 .1980 21.85 .4235 .3117 17.30 .5265 .3381 19.35 .5765 .3084 19.84 .5336 .3075
RegNeRF [6] (No kernel) 16.32 .3441 .5876 23.25 .7155 .2663 16.21 .2040 .6451 17.35 .4915 .4509 18.81 .5447 .4020 18.39 .4600 .4704
RegNeRF [6] (w/DP-kernel) 16.71 .3497 .4683 21.37 .6471 .1802 15.17 .1740 .6666 10.27 .1165 .6911 18.66 .5412 .4067 16.44 .3657 .4826
Sparse-DeRF (w/DN-kernel) - Ours 16.91 .3693 .5357 25.79 .7872 .1801 25.48 .6042 .2405 16.58 .4705 .3793 18.10 .5514 .3414 20.57 .5565 .3354
Sparse-DeRF (w/DP-kernel) - Ours 18.99 .4770 .4034 26.51 .8051 .1627 23.33 .4888 .2733 17.51 .5371 .3356 18.92 .5799 .3126 21.05 .5776 .2975
Real Scene Ball Basket Buick Coffee Decoration
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 21.61 .5043 .4992 17.23 fv.4384 .4746 19.20 .5508 .3579 23.21 .7341 .3281 14.78 .3299 .5640
MPR [56]+NeRF 21.46 .5046 .4992 16.98 .4260 .4766 19.39 .5486 .3629 22.98 .7267 .3284 14.88 .3460 .5532
Deblur-NeRF [1] 21.90 .5182 .4609 17.15 .4331 .4316 19.34 .5278 .3407 23.10 .7302 .2873 14.94 .3201 .5471
DP-NeRF [2] 23.20 .5842 .3847 17.42 .4046 .4658 19.26 .5323 .3374 24.05 .7597 .2624 14.98 .3018 .5618
RegNeRF [6] (No kernel) 21.16 .5100 .4952 18.12 .4711 .4569 20.11 .5821 .3412 26.14 .7837 .2965 11.35 .1045 .7030
RegNeRF [6] (w/DP-kernel) 19.13 .4391 .4042 17.42 .4165 .3746 13.31 .2697 .4715 16.40 .4684 .5239 11.07 .1058 .7089
Sparse-DeRF (w/DN-kernel) - Ours 22.28 .5506 .4334 18.69 .5058 .3774 19.28 .5414 .3371 26.76 .8126 .2419 17.12 .4192 .4761
Sparse-DeRF (w/DP-kernel) - Ours 23.39 .6010 .3806 20.41 .5451 .3305 19.48 .5531 .3251 27.77 .8364 .2049 16.33 .3914 .4950
Real Scene Girl Heron Parterre Puppet Stair Average
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 15.91 .5659 .4161 18.83 .4263 .4086 20.64 .4857 .4822 18.15 .4619 .4470 20.27 .3622 .5033 18.98 .4860 .4481
MPR [56]+NeRF 15.60 .5617 .4224 18.91 .4217 .4173 20.76 .4893 .4781 17.96 .4548 .4468 19.94 .3493 .4993 18.89 .4829 .4484
Deblur-NeRF [1] 15.62 .5351 .4206 18.85 .4350 .3398 19.71 .4318 .4549 17.77 .4400 .4358 20.49 .3901 .4326 18.89 .4761 .4151
DP-NeRF [2] 15.40 .5237 .4318 18.68 .4292 .3352 17.62 .2987 .4991 17.61 .4369 .4313 19.44 .3107 .4657 18.77 .4582 .4175
RegNeRF [6] (No kernel) 10.24 .1511 .7199 18.86 .4235 .4194 18.68 .4238 .5163 18.66 .4788 .4327 21.06 .3975 .4706 18.44 .4326 .4852
RegNeRF [6] (w/DP-kernel) 9.52 .1304 .7119 11.56 .0595 .6149 13.61 .1721 .7245 11.10 .1342 .6914 12.73 .0249 .6610 13.59 .2221 .5887
Sparse-DeRF (w/DN-kernel) - Ours 16.95 .5975 .3823 19.04 .4558 .3315 20.74 .4916 .4323 18.11 .4616 .4202 20.80 .3948 .4386 19.98 .5231 .3871
Sparse-DeRF (w/DP-kernel) - Ours 16.47 .5855 .3835 19.09 .4624 .3194 18.36 .3375 .4715 17.97 .4612 .4206 21.25 .4045 .4044 20.05 .5178 .3736
TABLE IX: Quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 6-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 17.12 .3663 .5663 23.12 .7160 .2713 28.42 .7463 .2106 19.92 .5889 .3866 19.68 .5749 .3853 21.65 .5985 .3638
MPR [56]+NeRF 17.21 .3754 .5649 23.06 .7087 .2749 28.38 .7461 .2093 20.23 .5966 .3783 19.71 .5728 .3870 21.72 .5999 .3629
Deblur-NeRF [1] 18.77 .5048 .3890 26.67 .8214 .1475 27.50 .7116 .1783 21.10 .6693 .2517 21.58 .6921 .2263 23.12 .6798 .2386
DP-NeRF [2] 21.63 .6402 .2984 27.63 .8475 .1224 25.36 .6227 .1861 21.27 .6818 .2023 22.49 .7259 .1899 23.68 .7036 .1998
RegNeRF [6] (No kernel) 17.04 .3690 .5729 23.40 .7205 .2649 18.05 .3610 .4870 20.33 .5980 .3745 19.63 .5760 .3833 19.69 .5249 .4165
RegNeRF [6] (w/DP-kernel) 21.03 .6273 .3076 27.73 .8455 .1238 18.04 .3173 .4987 21.84 .7246 .2022 19.36 .5661 .3908 21.60 .6162 .3046
Sparse-DeRF (w/DN-kernel) - Ours 18.33 .5144 .4004 26.63 .8184 .1513 28.16 .7344 .1755 22.06 .6988 .2245 21.43 .6855 .2380 23.32 .6903 .2379
Sparse-DeRF (w/DP-kernel) - Ours 21.29 .6179 .3261 27.34 .8340 .1298 28.19 .7365 .1606 22.30 .7233 .2017 22.21 .7159 .2036 24.27 .7255 .2044
Real Scene Ball Basket Buick Coffee Decoration
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 22.12 .5359 .4818 21.39 .6253 .3328 21.21 .6197 .3100 22.73 .7163 .3397 19.14 .5118 .4460
MPR [56]+NeRF 22.15 .5365 .4745 21.39 .6261 .3294 21.11 .6192 .3063 22.56 .7087 .3435 18.31 .4837 .4735
Deblur-NeRF [1] 24.47 .6473 .3342 23.63 .7204 .2108 21.45 .6285 .2625 23.70 .7604 .2588 17.74 .4790 .4296
DP-NeRF [2] 24.73 .6651 .3107 23.24 .6643 .2494 21.65 .6418 .2445 25.51 .7949 .2170 17.48 .4725 .4457
RegNeRF [6] (No kernel) 21.98 .5312 .4677 21.67 .6129 .3406 21.16 .6215 .3051 25.67 .7716 .3105 12.45 .1569 .7061
RegNeRF [6] (w/DP-kernel) 25.02 .6719 .2953 22.95 .6554 .2501 21.57 .6456 .2462 23.38 .7393 .2341 11.43 .1247 .6932
Sparse-DeRF (w/DN-kernel) - Ours 23.76 .6184 .3607 23.39 .7155 .2329 21.66 .6405 .2585 27.84 .8388 .2068 19.98 .5553 .3644
Sparse-DeRF (w/DP-kernel) - Ours 24.80 .6643 .3200 23.32 .6789 .2455 21.27 .6234 .2662 28.91 .8525 .1902 19.69 .5440 .3672
Real Scene Girl Heron Parterre Puppet Stair Average
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
Naive NeRF [19] 18.43 .6500 .3335 18.91 .4204 .4307 21.82 .5487 .4351 19.52 .5242 .3933 21.01 .3602 .5111 20.63 .5513 .4014
MPR [56]+NeRF 18.61 .6580 .3324 18.97 .4220 .4287 21.78 .5477 .4358 19.84 .5357 .3793 21.27 .3758 .4070 20.60 .5513 .4010
Deblur-NeRF [1] 18.09 .6582 .3055 19.40 .4709 .3166 21.92 .5557 .3872 20.58 .5834 .3226 22.60 .4993 .3349 21.36 .6003 .3163
DP-NeRF [2] 18.16 .6622 .3020 19.58 .4991 .2905 22.36 .5879 .3350 20.65 .5868 .2988 23.40 .5624 .2986 21.68 .6137 .2992
RegNeRF [6] (No kernel) 11.06 .1966 .7091 18.96 .4218 .4292 21.83 .5526 .4353 20.07 .5309 .3827 21.69 .3938 .4963 19.65 .4790 .4583
RegNeRF [6] (w/DP-kernel) 10.38 .1434 .7200 12.83 .1019 .5757 21.34 .5159 .3415 20.49 .5799 .3113 13.08 .2613 .6581 18.25 .4439 .4326
Sparse-DeRF (w/DN-kernel) - Ours 18.86 .6860 .2769 19.63 .5011 .2992 22.15 .5703 .3787 20.81 .5882 .3149 23.42 .5340 .3374 22.15 .6248 .3030
Sparse-DeRF (w/DP-kernel) - Ours 18.96 .6887 .2744 19.59 .5051 .2808 22.57 .5994 .3317 20.60 .5756 .3124 23.48 .5509 .3189 22.32 .6283 .2907
TABLE X: Ablation quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 2-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
Kernel Type Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
SS MGS PD PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DP-kernel 14.14 .2091 .6540 20.71 .5752 .3487 18.38 .3276 .4127 11.21 .1482 .6247 10.88 .1536 .6543 15.06 .2827 .5389
14.09 .2008 .6550 19.57 .5321 .3544 18.49 .3178 .4271 11.21 .1552 .6243 11.04 .1801 .6331 14.88 .2772 .5388
14.33 .2088 .6496 18.89 .5032 .3577 20.30 .3467 .4017 11.69 .1748 .6067 11.78 .1948 .6289 15.40 .2857 .5289
14.19 .2035 .6620 19.82 .5404 .3451 18.96 .3350 .4125 11.16 .1457 .6408 11.02 .1733 .6350 15.03 .2796 .5391
13.95 .1963 .6629 18.81 .5010 .3637 18.57 .3404 .4148 11.00 .1501 .6259 10.25 .1357 .6730 14.52 .2647 .5481
14.24 .2228 .6439 20.16 .5560 .3391 20.37 .3526 .4004 11.55 .1710 .5999 11.16 .1744 .6241 15.50 .2954 .5215
14.10 .2116 .6413 18.97 .5206 .3417 20.32 .3488 .4056 12.00 .1943 .5902 11.36 .1769 .6422 15.35 .2904 .5242
DN-kernel 14.14 .2009 .6647 20.58 .5768 .3443 18.37 .3247 .4061 11.56 .1704 .6075 11.05 .1690 .6425 15.14 .2884 .5330
14.21 .2177 .6650 20.33 .5663 .3591 18.74 .3200 .4093 11.13 .1547 .6191 10.78 .1554 .6566 15.04 .2828 .5418
14.20 .2107 .6686 21.02 .5999 .3326 19.64 .3330 .3950 11.55 .1804 .5903 11.48 .1883 .6351 15.12 .3025 .5243
14.11 .2096 .6588 18.78 .5094 .3928 19.76 .3464 .3901 10.58 .1206 .6565 10.99 .1648 .6289 14.84 .2702 .5454
14.24 .$2$6619 21.02 .5918 .3544 18.77 .3197 .4115 11.12 .1642 .6217 10.41 .1467 .6452 15.11 .2889 .5389
14.14 .2189 .6622 20.10 .$5$3761 20.11 .3372 .3939 11.49 .1753 .5975 11.99 .2026 .6299 15.57 .2979 .5319
14.27 .2200 .6564 20.57 .5675 .3589 19.81 .3330 .4058 11.67 .1827 .5947 11.28 .1800 .6298 15.52 .2966 .5291
Real Scene Ball Basket Buick Coffee Decoration
SS MGS PD PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DP-kernel 18.22 .4200 .5611 13.24 .2263 .6062 13.43 .2480 .5571 18.15 .5191 .4876 12.34 .1672 .6675
18.76 .4329 .5521 12.68 .2189 .6301 13.05 .2269 .5657 17.51 .4792 .4907 12.28 .1638 .6619
19.80 .4544 .5273 14.08 .2706 .5509 13.96 .3028 .4992 19.72 .6115 .4035 13.14 .2198 .6201
18.28 .4191 .5591 12.95 .2281 .5951 13.49 .2654 .5464 17.92 .5030 .4875 12.04 .1491 .6700
18.50 .4341 .5537 12.92 .2407 .5938 13.64 .2564 .5427 18.09 .5263 .4689 12.42 .1669 .6689
19.80 .4412 .5346 14.28 .2913 .5520 13.82 .2920 .5013 18.81 .5712 .4363 13.04 .2166 .6203
20.09 .4583 .5244 14.05 .2890 .5555 13.78 .2864 .5036 19.67 .6084 .4096 13.11 .2192 .6192
DN-kernel 18.80 .4325 .5521 12.86 .2414 .5947 13.11 .2446 .5609 18.66 .5272 .4696 12.09 .1508 .6722
18.88 .4388 .5519 13.03 .2516 .6018 13.36 .2533 .5523 18.04 .5145 .4847 12.38 .1736 .6545
19.95 .4596 .5394 13.52 .2534 .5871 14.11 .3087 .5051 19.33 .6003 .4133 13.13 .2153 .6222
18.70 .4235 .5673 13.24 .2505 .5931 13.52 .2520 .5438 18.15 .5031 .4949 12.32 .1683 .6617
18.36 .4140 .5667 13.14 .2482 .5975 13.40 .2549 .5461 17.90 .5098 .4857 12.37 .1634 .6658
20.06 .4637 .5299 13.56 .2593 .5828 14.51 .3204 .4869 19.40 .5956 .4198 13.07 .2189 .6230
20.07 .4612 .5290 13.62 .2678 .5835 14.38 .3182 .4913 19.51 .6048 .4146 13.09 .2173 .6256
Real Scene Girl Heron Parterre Puppet Stair Average
SS MGS PD PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DP-kernel 11.14 .2151 .6485 13.63 .1473 .5376 14.70 .2226 .6348 14.32 .2887 .5930 14.38 .0521 .6104 14.36 .2506 .5904
10.75 .1816 .6600 14.05 .1553 .5467 15.17 .2340 .6205 14.13 .2770 .5934 14.24 .0442 .6120 14.26 .2414 .5933
12.89 .3465 .5645 14.48 .1677 .5209 16.20 .2511 .5927 14.48 .2941 .5819 15.81 .1167 .6039 15.46 .3035 .5465
11.09 .2039 .6398 13.52 .1381 .5553 14.73 .2355 .6403 14.37 .2918 .5857 14.49 .0555 .6165 14.28 .2490 .5896
11.19 .2311 .6275 13.86 .1524 .5365 15.44 .2498 .6124 14.03 .2771 .6020 14.40 .0518 .6231 14.00 .2587 .5830
12.78 .3357 .5685 14.45 .1741 .5228 16.36 .2643 .5939 14.56 .2958 .5918 16.11 .1267 .6054 15.40 .3009 .5527
13.11 .3707 .5486 14.17 .1669 .5285 16.19 .2527 .5908 15.25 .3277 .5739 16.30 .1343 .6133 15.57 .3114 .5467
DN-kernel 10.94 .2031 .6436 13.97 .1459 .5499 15.02 .2320 .6431 14.28 .2764 .6036 14.40 .0525 .6313 14.41 .2506 .5921
10.43 .1696 .6654 13.10 .1336 .5652 15.09 .2075 .6536 14.66 .3000 .5950 14.36 .0528 .6122 14.33 .2495 .5937
12.55 .3184 .5736 14.51 .1850 .5217 16.65 .2830 .5953 14.68 .2944 .5924 15.93 .1189 .6081 15.44 .3037 .5558
11.19 .2214 .6475 13.95 .1421 .5508 14.30 .2189 .6739 14.99 .3117 .5894 14.35 .0574 .6164 14.47 .2549 .5939
11.55 .2299 .6195 13.93 .1543 .5382 14.92 .2291 .6399 14.74 .3042 .6001 14.43 .0556 .6211 14.47 .2563 .5881
12.70 .3438 .5627 14.57 .1779 .5200 16.51 .2872 .5825 14.24 .2848 .6097 16.02 .1216 .6018 15.46 .3073 .5519
12.85 .3418 .5585 14.30 .1706 .5197 16.61 .2903 .5912 14.91 .3141 .5886 15.97 .1255 .6125 15.53 .3112 .5515
TABLE XI: Ablation quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 4-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
Kernel Type Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
SS MGS PD PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DP-kernel 19.20 .5175 .3813 21.50 .6242 .1980 21.85 .4235 .3117 17.30 .5265 .3381 19.35 .5765 .3084 19.84 .5336 .3075
19.01 .4914 .4082 21.69 .6356 .1889 22.62 .4520 .3013 16.61 .4789 .3771 17.82 .5385 .3475 19.55 .5193 .3246
18.22 .4638 .4144 22.48 .6603 .1762 23.11 .4709 .2791 18.18 .5517 .3124 18.83 .5657 .3299 20.16 .5425 .3024
18.79 .4972 .3974 22.59 .6760 .2314 22.41 .4348 .3011 16.88 .4913 .3606 17.59 .5194 .3652 19.65 .5237 .3311
19.09 .5253 .3889 25.83 .7796 .1709 24.13 .5229 .2474 16.95 .5017 .3501 18.56 .5764 .3167 20.91 .5812 .2948
17.62 .4227 .4615 23.43 .7266 .1688 22.71 .4454 .2850 15.64 .4150 .4308 18.84 .5722 .3209 19.65 .5164 .3334
18.99 .4770 .4034 26.51 .8051 .1627 23.33 .4888 .2733 17.51 .5371 .3356 18.92 .5799 .3126 21.05 .5776 .2975
DN-kernel 17.26 .3740 .5088 25.51 .7767 .1897 23.38 .5068 .2649 15.91 .4209 .4198 17.91 .5213 .3661 19.99 .5199 .3499
16.89 .3611 .5353 25.82 .7876 .1770 22.92 .4820 .2809 15.52 .4381 .4173 18.07 .5317 .3684 19.89 .5201 .3558
16.33 .3387 .5636 25.55 .7813 .1874 24.28 .5544 .2649 16.60 .4635 .3781 17.96 .5095 .3870 19.44 .5295 .3562
16.49 .3444 .5691 25.40 .7674 .1982 21.87 .4194 .3203 16.48 .4798 .3830 18.38 .5412 .3445 19.72 .5104 .3630
17.11 .3766 .5195 25.43 .7770 .1891 22.41 .4585 .2864 16.63 .4792 .3826 18.57 .5559 .3476 20.03 .5294 .3450
16.22 .3389 .5637 25.12 .7566 .1950 23.00 .4792 .2810 17.37 .5088 .3661 18.08 .5275 .3677 19.96 .5222 .3547
16.91 .3693 .5357 25.79 .7872 .1801 25.48 .6042 .2405 16.58 .4705 .3793 18.10 .5514 .3414 20.57 .5565 .3354
Real Scene Ball Basket Buick Coffee Decoration
SS MGS PD PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DP-kernel 23.20 .5842 .3847 17.42 .4046 .4658 19.26 .5323 .3374 24.05 .7597 .2624 14.98 .3018 .5618
23.55 .6053 .3685 18.32 .4442 .4292 19.35 .5340 .3339 23.17 .7442 .2670 16.14 .3739 .5082
22.52 .5522 .3689 20.31 .5388 .3364 19.16 .5203 .3430 25.62 .7884 .2082 17.33 .4081 .4777
23.18 .5817 .3878 17.63 .4194 .4475 19.53 .5441 .3325 24.85 .7885 .2467 14.99 .3407 .5267
23.30 .5893 .3904 18.49 .4615 .3926 19.14 .5385 .3408 24.05 .7716 .2494 15.25 .3458 .5274
22.36 .5504 .3612 19.61 .5134 .3661 19.27 .5458 .3316 27.55 .8167 .2242 16.53 .3902 .4867
23.39 .6010 .3806 20.41 .5451 .3305 19.48 .5531 .3251 27.77 .8364 .2049 16.33 .3914 .4950
DN-kernel 21.90 .5182 .4609 17.15 .4331 .4316 19.34 .5278 .3407 23.10 .7302 .2873 14.94 .3201 .5471
21.94 .5220 .4571 17.38 .4464 .4298 19.26 .5401 .3418 23.03 .7264 .2830 14.82 .3293 .5363
21.85 .5168 .4514 17.87 .4641 .4014 19.23 .5289 .3467 26.91 .8082 .2473 16.69 .4031 .4859
21.85 .5313 .4392 16.97 .4301 .4228 19.45 .5443 .3403 23.78 .7679 .2754 15.18 .3549 .5112
22.06 .5319 .4481 16.75 .4308 .4333 19.34 .5421 .3373 23.44 .7421 .2943 14.86 .3429 .5299
22.09 .5299 .4477 17.81 .4854 .3812 19.31 .5400 .3384 27.40 .8263 .2353 16.93 .4023 .4879
22.28 .5506 .4334 18.69 .5058 .3774 19.28 .5414 .3371 26.76 .8126 .2419 17.12 .4192 .4761
Real Scene Girl Heron Parterre Puppet Stair Average
SS MGS PD PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DP-kernel 15.40 .5237 .4318 18.68 .4292 .3352 17.62 .2987 .4991 17.61 .4369 .4313 19.44 .3107 .4657 18.77 .4582 .4175
15.67 .5342 .4274 18.44 .4145 .3322 17.37 .2900 .4935 18.49 .4765 .3982 19.85 .3433 .4503 19.04 .4760 .4008
17.03 .6045 .3636 18.90 .4526 .3185 18.38 .3323 .4575 17.89 .4408 .4333 20.33 .3416 .4297 19.75 .4980 .3737
14.80 .4950 .4538 18.83 .4336 .3308 17.81 .3089 .4915 18.12 .4505 .4205 20.23 .3615 .4422 19.00 .4724 .4080
15.57 .5297 .4349 19.04 .4578 .3278 17.88 .3094 .4887 17.84 .4461 .4329 19.47 .3102 .4522 19.00 .4760 .4037
16.78 .5919 .3806 18.99 .4575 .3166 18.28 .$3$4630 17.63 .4272 .4243 20.36 .3594 .4201 19.74 .4986 .3774
16.47 .5855 .3835 19.09 .4624 .3194 18.36 .3375 .4715 17.97 .4612 .4206 21.25 .4045 .4044 20.05 .5178 .3736
DN-kernel 15.62 .5351 .4206 18.85 .4350 .3398 19.71 .4318 .4549 17.77 .4400 .4358 20.49 .3901 .4326 18.89 .4761 .4151
15.04 .5147 .4312 18.75 .4291 .3585 20.73 .4841 .4379 18.15 .4647 .4248 20.45 .3825 .4249 18.96 .4839 .4125
17.13 .6103 .3623 18.96 .4527 .3366 20.47 .4735 .4464 18.39 .4694 .4156 20.94 .3914 .4432 19.84 .5118 .3937
15.17 .5215 .4383 18.97 .4367 .3480 20.57 .4833 .4463 18.24 .4620 .4252 20.42 .3752 .4578 19.06 .4907 .4105
15.48 .5272 .4282 19.00 .4430 .3458 20.37 .4801 .4401 18.08 .4577 .4276 20.05 .3635 .4601 18.94 .4861 .4145
17.07 .6081 .3706 18.95 .4514 .3254 20.64 .4854 .4385 18.03 .4510 .4298 21.05 .4086 .4182 19.93 .5188 .3873
16.95 .5975 .3823 19.04 .4558 .3315 20.74 .4916 .4323 18.11 .4616 .4202 20.80 .3948 .4386 19.98 .5231 .3871
TABLE XII: Ablation quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 6-view settings. Each color shading represents the best, second best and third best result, respectively. DP-kernel and DN-kernel denote the kernel of [2] and [1].
Kernel Type Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
SS MGS PD PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DP-kernel 21.63 .6402 .2984 27.63 .8475 .1224 25.36 .6227 .1861 21.27 .6818 .2023 22.49 .7259 .1899 23.68 .7036 .1998
21.84 .6556 .2947 27.48 .8419 .1272 26.73 .6891 .1687 22.57 .7289 .1950 22.15 .7160 .2015 24.15 .7263 .1974
19.30 .5628 .3553 27.56 .8408 .1262 28.10 .7374 .1544 21.83 .7128 .2074 22.34 .7197 .1928 23.83 .7147 .2072
20.52 .5994 .3224 27.29 .8322 .1349 27.07 .7002 .1719 22.50 .7306 .1958 22.07 .7127 .2011 23.89 .7150 .2052
21.27 .6104 .3191 27.39 .8376 .1301 28.54 .7518 .1608 22.32 .7290 .1966 21.87 .7109 .2037 24.28 .7279 .2021
18.43 .5407 .3738 27.44 .8399 .1313 27.89 .7321 .1596 22.02 .7197 .2071 21.87 .7118 .2046 23.53 .7088 .2153
21.29 .6179 .3261 27.34 .8340 .1298 28.19 .7365 .1606 22.30 .7233 .2017 22.21 .7159 .2036 24.27 .7255 .2044
DN-kernel 18.77 .5048 .3890 26.67 .8214 .1475 27.50 .7116 .1783 21.10 .6693 .2517 21.58 .6921 .2263 23.12 .6798 .2386
17.87 .4554 .4370 26.76 .8239 .1462 28.12 .7316 .1735 21.69 .6979 .2282 21.61 .6887 .2294 23.59 .6795 .2429
18.06 .4849 .4104 26.54 .8159 .1486 28.09 .7392 .1725 20.97 .6866 .2394 21.71 .6895 .2320 24.33 .6832 .2406
19.22 .5397 .3756 26.52 .8184 .1514 27.72 .7242 .1828 21.67 .6961 .2306 21.64 .6920 .2300 23.35 .6941 .2341
19.77 .5461 .3692 26.63 .8192 .1480 28.17 .7353 .1733 21.08 .6878 .2367 21.39 .6821 .2347 23.41 .6941 .2324
17.50 .4349 .4613 26.72 .8226 .1491 27.40 .7046 .1848 21.73 .6962 .2369 21.20 .6688 .2477 22.91 .6654 .2560
18.33 .5144 .4004 26.63 .8184 .1513 28.16 .7344 .1755 22.06 .6988 .2245 21.43 .6855 .2380 23.32 .6903 .2379
Real Scene Ball Basket Buick Coffee Decoration
SS MGS PD PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DP-kernel 24.73 .6651 .3107 23.24 .6643 .2494 21.65 .6418 .2445 25.52 .7949 .2170 17.48 .4725 .4457
24.93 .6697 .3054 22.91 .6688 .2453 21.48 .6400 .2580 28.12 .8224 .1987 18.34 .4934 .4226
24.21 .6412 .3094 22.82 .6631 .2508 21.17 .6251 .2681 28.39 .8448 .1722 19.82 .5485 .3697
24.92 .6661 .3111 22.27 .6326 .2860 21.30 .6234 .2710 26.62 .7973 .2259 18.07 .4926 .4230
24.86 .6598 .3181 21.29 .6094 .3065 21.40 .6359 .2584 28.58 .8332 .1910 17.78 .4869 .4352
25.01 .6691 .3007 23.19 .6706 .2481 21.12 .6212 .2662 28.23 .8410 .1798 19.88 .5550 .3606
24.80 .6643 .3200 23.32 .6789 .2455 21.27 .6234 .2662 28.91 .8525 .1902 19.69 .5440 .3672
DN-kernel 24.47 .6473 .3342 23.63 .7204 .2108 21.45 .6285 .2625 23.70 .7604 .2588 17.74 .4790 .4296
23.02 .5757 .3964 22.96 .7128 .2230 21.72 .6389 .2595 26.90 .8004 .2354 17.21 .4774 .4346
24.15 .6343 .3465 22.59 .6894 .2365 21.30 .6185 .2678 26.41 .7947 .2300 19.38 .5272 .3908
23.68 .6175 .3572 22.90 .7035 .2255 20.95 .6172 .2716 27.18 .8083 2336 18.53 .5030 .4084
24.60 .6530 .3439 22.96 .7132 .2296 21.43 .6360 .2642 26.13 .8178 .2159 18.07 .4784 .4420
23.93 .6332 .3464 23.89 .7342 .2188 21.11 .6247 .2747 26.88 .8074 .2403 19.53 .5383 .3855
23.76 .6184 .3607 23.39 .7155 .2329 21.66 .6405 .2585 27.84 .8388 .2068 19.98 .5553 .3644
Real Scene Girl Heron Parterre Puppet Stair Average
SS MGS PD PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DP-kernel 18.16 .6622 .3020 19.58 .4991 .2905 22.36 .5879 .3350 20.65 .5868 .2988 23.40 .5624 .2986 21.68 .6137 .2992
18.79 .6771 .2890 19.56 .4967 .2903 22.53 .5972 .3318 20.41 .5765 .3101 23.96 .5723 .2997 22.10 .6214 .2951
18.77 .6833 .2703 19.58 .4998 .2856 22.27 .5801 .3440 20.48 .5790 .3043 23.23 .5468 .3149 22.07 .6212 .2889
18.50 .6650 .2956 19.59 .4895 .3003 22.45 .5926 .3344 20.29 .5687 .3230 23.33 .5545 .3120 21.73 .6082 .3082
18.36 .6628 .2989 19.57 .4920 .3006 22.69 .6015 .3296 20.38 .5774 .3211 23.47 .5462 .3244 21.84 .6105 .3084
19.20 .6872 .2695 19.45 .4947 .2871 22.33 .5833 .3438 20.46 .5800 .3035 23.34 .5550 .3082 $2$6257 .2868
18.96 .6887 .2744 19.59 .5051 .2808 22.57 .5994 .3317 20.60 .5756 .3124 23.48 .5509 .3189 22.32 .6283 .2907
DN-kernel 18.09 .6582 .3055 19.40 .4709 .3166 21.92 .5557 .3872 20.58 .5834 .3226 22.60 .4993 .3349 21.36 .6003 .3163
18.19 .6615 .3009 19.42 .4879 .3052 22.18 .5758 .3733 20.49 .5823 .3167 23.18 .5312 .3330 21.53 .6044 .3178
18.93 .6856 .2748 19.58 .4904 .3122 21.91 .5570 .3853 20.44 .5810 .3232 22.90 .5231 .3294 21.76 .6101 .3097
18.18 .6538 .3005 19.65 .4962 .3083 22.34 .5834 .3608 20.41 .5785 .3305 23.08 .5257 .3458 21.69 .6087 .3142
18.74 .6782 .2855 19.74 .4948 .3153 22.25 .5796 .3658 20.45 .5771 .3183 22.63 .5009 .3637 21.70 .6129 .3144
18.53 .6646 .2923 19.45 .4859 .3182 21.62 .5510 .3909 20.71 .5819 .3126 23.18 .5332 .3325 21.88 .6154 .3112
18.86 .6860 .2769 19.63 .5011 .2992 22.15 .5703 .3787 20.81 .5882 .3149 23.42 .5340 .3374 22.15 .6248 .3030
TABLE XIII: Quantitative results of the RegNeRF [6] with and without the blur kernel from 2-view to 10-view settings. For comparison, the blur kernel of the DP-NeRF [2] is employed, which is denoted as DP-kernel. Color shading represents the better result.
Decoration Scene # of views
PSNR($\uparrow$) Blur Kernel 2 3 4 5 6 7 8 9 10
RegNeRF [6] $\times$ 11.20 11.28 11.43 12.11 12.45 14.61 16.69 18.41 20.21
RegNeRF [6] DP-kernel 11.15 10.84 11.07 11.54 11.43 13.50 16.16 18.56 21.75
SSIM($\uparrow$) Blur Kernel 2 3 4 5 6 7 8 9 10
RegNeRF [6] $\times$ .1064 .1262 .1139 .1589 .1569 .2876 .3956 .4880 .5508
RegNeRF [6] DP-kernel .0927 .1115 .1058 .1228 .1247 .2034 .3645 .4986 .6526
LPIPS($\downarrow$) Blur Kernel 2 3 4 5 6 7 8 9 10
RegNeRF [6] $\times$ .7475 .7059 .6995 .7078 .7061 .6258 .5588 .4818 .4322
RegNeRF [6] DP-kernel .7520 .7191 .7089 .6992 .6932 .6547 .5277 .4178 .2804
TABLE XIV: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 2-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
Synthetic Scene Blur Kernel Factory Cozyroom Pool Tanabata Trolley Average
Model PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 14.57 .2523 .6680 17.13 .4616 .3808 14.23 .1473 .7298 13.01 .2408 .5956 14.04 .3227 .5602 14.60 .2849 .5869
RegNeRF [6] DP-kernel 14.20 .2214 .6412 18.88 .5136 .3476 13.03 .1003 .6845 9.21 .0815 .6920 10.86 .1644 .6657 13.24 .2162 .6062
Sparse-DeRF(Ours) DP-kernel 14.10 .2116 .6413 18.97 .5206 .3417 20.32 .3488 .4056 12.00 .1943 .5902 11.36 .1769 .6422 15.35 .2904 .5242
Real Scene Blur Kernel Ball Basket Buick Coffee Decoration
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 20.48 .4951 .5398 15.53 .3388 .5403 17.05 .4241 .4361 23.24 .6950 .3505 10.94 .0816 .7545
RegNeRF [6] DP-kernel 18.58 .4297 .5565 13.49 .2429 .5985 12.51 .2199 .5579 15.75 .4280 .5370 11.15 .0927 .7520
Sparse-DeRF(Ours) DP-kernel 20.09 .4583 .5244 14.05 .2890 .5555 13.78 .2864 .5036 19.67 .6084 .4096 13.11 .2192 .6192
Real Scene Blur Kernel Girl Heron Parterre Puppet Stair Average
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 8.31 .0602 .7545 11.43 .0951 .6687 15.58 .2709 .6510 16.02 .3548 .5739 16.30 .1816 .6182 15.49 .2997 .5888
RegNeRF [6] DP-kernel 7.86 .0419 .7737 11.57 .0783 .6246 12.54 .1524 .7091 10.67 .1331 .6604 13.49 .0170 .6770 12.76 .1836 .6447
Sparse-DeRF(Ours) DP-kernel 13.11 .3707 .5486 14.17 .1669 .5285 16.19 .2527 .5908 15.25 .3277 .5739 16.30 .1343 .6133 15.57 .3114 .5467
TABLE XV: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 4-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
Synthetic Scene Blur Kernel Factory Cozyroom Pool Tanabata Trolley Average
Model PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 16.32 .3441 .5876 23.25 .7155 .2663 16.21 .2040 .6451 17.35 .4915 .4509 18.81 .5447 .4020 18.39 .4600 .4704
RegNeRF [6] DP-kernel 16.71 .3497 .4683 21.37 .6471 .1802 15.17 .1740 .6666 10.27 .1165 .6911 18.66 .5412 .4067 16.44 .3657 .4826
Sparse-DeRF(Ours) DP-kernel 18.99 .4770 .4034 26.51 .8051 .1627 23.33 .4888 .2733 17.51 .5371 .3356 18.92 .5799 .3126 21.05 .5776 .2975
Real Scene Blur Kernel Ball Basket Buick Coffee Decoration
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 21.16 .5100 .4952 18.12 .4711 .4569 20.11 .5821 .3412 26.14 .7837 .2965 11.35 .1045 .7030
RegNeRF [6] DP-kernel 19.13 .4391 .4042 17.42 .4165 .3746 13.31 .2697 .4715 16.40 .4684 .5239 11.07 .1058 .7089
Sparse-DeRF(Ours) DP-kernel 23.39 .6010 .3806 20.41 .5451 .3305 19.48 .5531 .3251 27.77 .8364 .2049 16.33 .3914 .4950
Real Scene Blur Kernel Girl Heron Parterre Puppet Stair Average
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 10.24 .1511 .7199 18.86 .4235 .4194 18.68 .4238 .5163 18.66 .4788 .4327 21.06 .3975 .4706 18.44 .4326 .4852
RegNeRF [6] DP-kernel 9.52 .1304 .7119 11.56 .0595 .6149 13.61 .1721 .7245 11.10 .1342 .6914 12.73 .0249 .6610 13.59 .2221 .5887
Sparse-DeRF(Ours) DP-kernel 16.47 .5855 .3835 19.09 .4624 .3194 18.36 .3375 .4715 17.97 .4612 .4206 21.25 .4045 .4044 20.05 .5178 .3736
TABLE XVI: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 6-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
Synthetic Scene Blur Kernel Factory Cozyroom Pool Tanabata Trolley Average
Model PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 17.04 .3690 .5729 23.40 .7205 .2649 18.05 .3610 .4870 20.33 .5980 .3745 19.63 .5760 .3833 19.69 .5249 .4165
RegNeRF [6] DP-kernel 21.03 .6273 .3076 27.73 .8455 .1238 18.04 .3173 .4987 21.84 .7246 .2022 19.36 .5661 .3908 21.60 .6162 .3046
Sparse-DeRF(Ours) DP-kernel 21.29 .6179 .3261 27.34 .8340 .1298 28.19 .7365 .1606 22.30 .7233 .2017 22.21 .7159 .2036 24.27 .7255 .2044
Real Scene Blur Kernel Ball Basket Buick Coffee Decoration
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 21.98 .5312 .4677 21.67 .6129 .3406 21.16 .6215 .3051 25.67 .7716 .3105 12.45 .1569 .7061
RegNeRF [6] DP-kernel 25.02 .6719 .2953 22.95 .6554 .2501 21.57 .6456 .2462 23.38 .7393 .2341 11.43 .1247 .6932
Sparse-DeRF(Ours) DP-kernel 24.80 .6643 .3200 23.32 .6789 .2455 21.27 .6234 .2662 28.91 .8525 .1902 19.69 .5440 .3672
Real Scene Blur Kernel Girl Heron Parterre Puppet Stair Average
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 11.06 .1966 .7091 18.96 .4218 .4292 21.83 .5526 .4353 20.07 .5309 .3827 21.69 .3938 .4963 19.65 .4790 .4583
RegNeRF [6] DP-kernel 10.38 .1434 .7200 12.83 .1019 .5757 21.34 .5159 .3415 20.49 .5799 .3113 13.08 .2613 .6581 18.25 .4439 .4326
Sparse-DeRF(Ours) DP-kernel 18.96 .6887 .2744 19.59 .5051 .2808 22.57 .5994 .3317 20.60 .5756 .3124 23.48 .5509 .3189 22.32 .6283 .2907
TABLE XVII: Quantitative comparison between naive gradient scaling of [51] and our MGS for entire scenes of synthetic and real scenes obtained from 2-view settings. The experimental results are presented according to the two types of kernel we utilize in this paper, which are kernels of the DP-NeRF [2] and Deblur-NeRF [1]. Each kernel is denoted as DP-kernel and DN-kernel, respectively. Color shading represents the better result.
Synthetic Scene Gradient Scaling Factory Cozyroom Pool Tanabata Trolley Average
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel $+$ Naive [51] 11.93 .1340 .6835 17.82 .4573 .4827 18.89 .3131 .4237 11.45 .1603 .6107 11.12 .1769 .6207 14.70 .2483 .5643
DN-kernel $+$ MGS 14.20 .2107 .6686 21.02 .5999 .3326 20.30 .3467 .4017 11.55 .1804 .5903 11.48 .1883 .6351 15.12 .3025 .5243
DP-kernel $+$ Naive [51] 11.98 .1378 .6799 16.54 .3988 .5360 19.08 .3135 .4275 11.30 .1653 .5992 11.14 .1731 .6291 14.01 .2377 .5743
DP-kernel $+$ MGS 14.33 .2088 .6496 18.89 .5032 .3577 19.64 .3330 .3950 11.69 .1748 .6067 11.78 .1948 .6289 15.40 .2857 .5289
Real Scene Gradient Scaling Ball Basket Buick Coffee Decoration
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel $+$ Naive [51] 17.96 .4198 .5715 13.27 .2431 .5864 13.88 .2982 .5075 19.17 .5960 .4223 12.08 .2076 .6294
DN-kernel $+$ MGS 19.95 .4596 .5394 13.52 .2534 .5871 14.11 .3087 .5051 19.33 .6003 .4133 13.13 .2153 .6222
DP-kernel $+$ Naive [51] 17.85 .4053 .5819 13.43 .2445 .5694 13.97 .3061 .5014 19.36 .6070 .4123 12.81 .2071 .6238
DP-kernel $+$ MGS 19.80 .4544 .5273 14.08 .2706 .5509 13.96 .3028 .4992 19.72 .6115 .4035 13.14 .2198 .6201
Real Scene Gradient Scaling Girl Heron Parterre Puppet Stair Average
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel $+$ Naive [51] 12.40 .2923 .5883 13.46 .1458 .5483 16.52 .2778 .5930 13.76 .2670 .5948 15.99 .1122 .6118 14.85 .2860 .5653
DN-kernel $+$ MGS 12.55 .3184 .5736 14.51 .1850 .5217 16.65 .2830 .5953 14.68 .2944 .5924 15.93 .1189 .6081 15.44 .3037 .5558
DP-kernel $+$ Naive [51] 12.84 .3408 .5621 14.32 .1672 .5234 16.22 .2547 .5963 12.85 .2309 .6045 15.80 .1072 .6159 14.95 .2871 .5591
DP-kernel $+$ MGS 12.89 .3465 .5645 14.48 .1677 .5209 16.20 .2511 .5927 14.48 .2941 .5819 15.81 .1167 .6039 15.46 .3035 .5465
TABLE XVIII: Quantitative comparison between naive gradient scaling of [51] and our MGS for entire scenes of synthetic and real scenes obtained from 4-view settings. The experimental results are presented according to the two types of kernel we utilize in this paper, which are the kernels of DP-NeRF [2] and Deblur-NeRF [1]. Each kernel is denoted as DP-kernel and DN-kernel, respectively. Color shading represents the better result.
Synthetic Scene Gradient Scaling Factory Cozyroom Pool Tanabata Trolley Average
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel $+$ Naive [51] 12.80 .1949 .6631 23.41 .7202 .2517 23.54 .5069 .2508 14.22 .3275 .5065 15.41 .4109 .4487 18.69 .4321 .4242
DN-kernel $+$ MGS 16.33 .3387 .5636 25.55 .7813 .1874 24.28 .5544 .2649 16.60 .4635 .3781 17.96 .5095 .3870 19.44 .5295 .3562
DP-kernel $+$ Naive [51] 11.89 .1392 .6831 22.26 .6676 .2354 22.36 .4721 .2616 14.14 .3024 .4953 15.62 .3850 .4351 17.25 .3933 .4221
DP-kernel $+$ MGS 18.22 .4638 .4144 22.48 .6603 .1762 23.11 .4709 .2791 18.18 .5517 .3124 18.83 .5657 .3299 20.16 .5425 .3024
Real Scene Gradient Scaling Ball Basket Buick Coffee Decoration
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel $+$ Naive [51] 21.01 .4947 .4699 16.89 .4274 .4394 18.79 .5123 .3654 26.36 .7986 .2604 16.77 .4036 .4896
DN-kernel $+$ MGS 21.85 .5168 .4514 17.87 .4641 .4014 19.23 .5289 .3467 26.91 .8082 .2473 16.69 .4031 .4859
DP-kernel $+$ Naive [51] 20.06 .4560 .4361 18.23 .4615 .3970 18.65 .5074 .3611 24.80 .7541 .2497 16.32 .3843 .4878
DP-kernel $+$ MGS 22.52 .5522 .3689 20.31 .5388 .3364 19.16 .5203 .3430 25.62 .7884 .2082 17.33 .4081 .4777
Real Scene Gradient Scaling Girl Heron Parterre Puppet Stair Average
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel $+$ Naive [51] 16.83 .5976 .3730 18.60 .4154 .3632 20.10 .4602 .4612 16.90 .4247 .4646 20.53 .3749 .4319 19.28 .4909 .4119
DN-kernel $+$ MGS 17.13 .6103 .3623 18.96 .4527 .3366 20.47 .4735 .4464 18.39 .4694 .4156 20.94 .3914 .4432 19.84 .5118 .3937
DP-kernel $+$ Naive [51] 16.31 .5787 .3874 18.85 .4481 .3152 18.21 .3274 .4750 16.74 .4101 .4634 20.32 .3421 .4415 18.85 .4670 .4014
DP-kernel $+$ MGS 17.03 .6045 .3636 18.90 .4526 .3185 18.38 .3323 .4575 17.89 .4408 .4333 20.33 .3416 .4297 19.75 .4980 .3737
TABLE XIX: Quantitative comparison between naive gradient scaling of [51] and our MGS for entire scenes of synthetic and real scenes obtained from 6-view settings. The experimental results are presented according to the two types of kernel we utilize in this paper, which are the kernels of DP-NeRF [2] and Deblur-NeRF [1]. Each kernel is denoted as DP-kernel and DN-kernel, respectively. Color shading represents the better result.
Synthetic Scene Gradient Scaling Factory Cozyroom Pool Tanabata Trolley Average
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel $+$ Naive [51] 14.70 .2560 .6091 26.13 .8080 .1658 27.24 .7167 .1810 20.29 .6379 .2816 18.48 .5824 .3208 22.00 .6002 .3117
DN-kernel $+$ MGS 18.06 .4849 .4104 26.54 .8159 .1486 28.09 .7392 .1725 20.97 .6866 .2394 21.71 .6895 .2320 22.40 .6832 .2406
DP-kernel $+$ Naive [51] 14.47 .2649 .5744 26.71 .8233 .1456 27.90 .7330 .1598 20.04 .6253 .2528 18.63 .6020 .2865 21.55 .6097 .2838
DP-kernel $+$ MGS 19.30 .5628 .3553 27.56 .8408 .1262 28.10 .7374 .1544 21.83 .7128 .2074 22.34 .7197 .1928 23.83 .7147 .2072
Real Scene Gradient Scaling Ball Basket Buick Coffee Decoration
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel $+$ Naive [51] 22.07 .5574 .4040 21.76 .6780 .2515 21.03 .6107 .2837 27.69 .8236 .2275 19.72 .5396 .3770
DN-kernel $+$ MGS 24.15 .6343 .3465 22.59 .6894 .2365 21.30 .6185 .2678 26.41 .7947 .2300 19.38 .5272 .3908
DP-kernel $+$ Naive [51] 24.04 .6385 .3293 21.29 .6075 .3275 20.07 .6133 .2724 28.42 .8342 .1931 19.60 .5366 .3727
DP-kernel $+$ MGS 24.21 .6412 .3094 22.82 .6631 .2508 21.17 .6251 .2681 28.39 .8448 .1722 19.82 .5485 .3697
Real Scene Gradient Scaling Girl Heron Parterre Puppet Stair Average
Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
DN-kernel $+$ Naive [51] 18.68 .6777 .2824 19.41 .4885 .3078 21.67 .5481 .3963 18.24 .5173 .3792 21.75 .4360 .3543 21.20 .5877 .3274
DN-kernel $+$ MGS 18.93 .6856 .2748 19.58 .4904 .3122 21.91 .5570 .3853 20.44 .5810 .3232 22.90 .5231 .3294 21.76 .6101 .3097
DP-kernel $+$ Naive [51] 18.70 .6770 .2789 19.13 .4822 .2918 21.97 .5665 .3548 18.36 .5114 .3739 23.27 .5432 .3167 21.49 .6010 .3111
DP-kernel $+$ MGS 18.77 .6833 .2703 19.58 .4998 .2856 22.27 .5801 .3440 20.48 .5790 .3043 23.23 .5468 .3149 22.07 .6212 .2889

### v2

@@
 Sparse-DeRF successfully regularizes the complicated joint optimization, presenting alleviated overfitting artifacts and enhanced quality on radiance fields.
-The regularization consists of three key components: Surface smoothness, helps the model accurately predict the scene structure utilizing unseen and additional hidden rays derived from the blur kernel based on statistical tendencies of real-world; Modulated gradient scaling, helps the model adjust the amount of the backpropagated gradient according to the arrangements of scene objects; Perceptual distillation improves the perceptual quality by overcoming the ill-posed multi-view inconsistency of image deblurring and distilling the pre-filtered information, compensating for the lack of clean information in blurry images.
+The regularization consists of three key components: Surface smoothness, helps the model accurately predict the scene structure utilizing unseen and additional hidden rays derived from the blur kernel based on statistical tendencies of real-world; Modulated gradient scaling, helps the model adjust the amount of the backpropagated gradient according to the arrangements of scene objects; Perceptual distillation improves the perceptual quality by overcoming the ill-posed multi-view inconsistency of image deblurring and distilling the pre-deblurred information, compensating for the lack of clean information in blurry images.
 We demonstrate the effectiveness of the Sparse-DeRF with extensive quantitative and qualitative experimental results by training DeRF from 2-view, 4-view, and 6-view blurry images.
@@
 To solve this problem, several NeRF studies [1, 2, 3] have attempted to construct deblurred neural radiance fields (hereafter, DeRF) from blurry images using joint optimization of internal implicit blur kernel and radiance fields, but they use dozens of blurry images to train, which is actually not practical scenarios.
-The assumed experimental environments, where radiance fields are trained from about 20 to 30 blurry images, seem unlikely to occur in reality Hence we delve into the practical consideration for situations where only blurry images are utilized.
+The assumed experimental environments, where radiance fields are trained from about 20 to 30 blurry images, seem unlikely to occur in reality.
+Hence we delve into the practical consideration for situations where only blurry images are utilized.
 We reasoned that situations requiring the use of only blurry images would arise when the available images for reconstructing the desired 3D space are both very limited and blurry.
 Following this rationale, we propose a novel pragmatic scenario for radiance fields from blurry images that establish the DeRF from sparse view settings.
-Specifically, we set the 2-view, 4-view, and 6-view settings based on our consideration of the practical applications of research on generating radiance fields from blurry images.
+Specifically, we set the 2, 4, and 6-view settings based on our consideration of the practical applications of research on generating radiance fields from blurry images.
 Actually, the NeRF system already has an inherent drawback: it is prone to be overfitted to training views and struggles to grasp correct geometry when only sparse view inputs are available.
@@
 Furthermore, in the DeRF system, it is challenging to use other data-deriven priors such as predicted depth supervision since available images do not ensure the confidence of the estimated values due to the inherent degradation of the given images.
-Therefore, our goal is to regularize the complex joint optimization of blur kernel and radiance fields for DeRF to enhance the structural and perceptual quality of radiance fields from sparse blurry images, overcoming the aforementioned challenging issues.
+The inherent degradations cause inconsistency across multiple views, which hinder accurate geometry prediction in NeRF, when it is used as additional prior.
+This tendency becomes more pronounced as the number of input images increases.
+We validated the tendency through experimental results.
+Therefore, our goal is to regularize the complex joint optimization of blur kernel and radiance fields for DeRF to enhance the structural and perceptual quality of radiance fields from sparse blurry images, overcoming the aforementioned challenging issues without pretrained depth prior.
 In this paper, we propose for the first time to ameliorate the spatial ambiguity and enhance the sharp texture of the DeRF from sparse view, which we refer to as Sparse-DeRF.
@@
 These two geometric constraints improve the structural scene geometry of radiance fields even without explicit depth supervision in a sparse view setting.
-In addition to geometric constraints, we propose the perceptual distillation (PD) as a perceptual prior to enhance the detailed texture of the radiance fields by taking advantage of the previously established image deblurring algorithm.
-Traditional image deblurring has shown significant performance improvements alongside the advancement of deep learning, demonstrating more enhanced details and textures.
-We believe that the sharp texture information from such deblurred images can be used as additional complementary information to achieve high fidelity in the Sparse-DeRF environment, where only a few degraded images are available to reconstruct the scene.
-However, while we can take the pre-filtered images with a pre-trained deep learning-based image deblurring model, the independence of image deblurring poses challenges in directly utilizing deblurred images as pixel-wise color supervision, due to inconsistency across the given images.
+In addition to geometric constraints, we introduce the perceptual distillation (PD) as a perceptual prior to enhance the detailed texture of the radiance fields by taking advantage of the established image deblurring algorithm.
+Traditional image deblurring has shown significant performance alongside the advancement of deep learning, demonstrating more enhanced details and textures.
+We believe that the sharp texture information from such deblurred images can be used as complementary information to achieve high fidelity in the Sparse-DeRF environment, where only a few degraded images are available to reconstruct the scene.
+However, while we can take the pre-deblurred images with a pre-trained deep learning-based image deblurring model, the independence of image deblurring poses challenges in directly utilizing deblurred images as pixel-wise color supervision, due to inconsistency across the given images.
 This inconsistency comes from the inherent ill-posed property of the image deblurring that breaks the geometric and appearance consistency across the multi-view images of the single 3D scene.
-Hence, we impart the perceptual information of pre-filtered images to the radiance fields by distilling the features extracted from the deep learning-based image feature extractor.
+Hence, we impart the perceptual information of pre-deblurred images to the radiance fields by distilling the features extracted from the deep learning-based image feature extractor.
 Extracted features enable the radiance fields to enhance perceptual quality by utilizing pre-deblurred textures.
@@
 In addition, its implicit representation capability leads to explosive development of other graphical tasks such as modeling dynamic scenes [28, 29, 30, 31], relighting [32, 33], 3D reconstructions [34, 35], and human avatar [36].
-II-B Radiance Fields in Practical Scenarios
+II-B Radiance Fields from Degraded Images
+Establishing NeRF from degraded images is recently emerging since the ideal training condition in images for NeRF often breaks in real-world scenarios.
+RawNeRF [37] denoises the internal noise of the camera sensor to construct high dynamic range (HDR) radiance fields from dark raw images and controls the camera exposure.
+Similarly, NaN [38] deals with burst noise in images, generating denoised images based on IBRNet [39], which is another image-based rendering approach.
+For more practical applications of NeRF in real-world, DeblurNeRF [1] firstly attempts to deal with two types of blur degradation in images, blur from camera motion and defocus, constructing deblurred neural radiance fields (DeRF) from only blurry images.
+They imitate the blurring process integrating the concept of blind deblurring in image deblurring with the NeRF system, modeling the blur kernel as pixel-wise independent ray transformation and composition weights to approximate the blurring process.
+Another representative approach is DP-NeRF [2], which is a key prior work that provided strong motivation for our current research and serves as the main baseline in our experiments.
+DP-NeRF [2] imposes physical consistency across the images by modeling the blur kernel as the 3D rigid transformation of rays depending on each view, to approximate the actual blurring process in the camera more precisely.
+Several approaches [40, 41, 3] are also proposed, attempting to improve the quality of the constructed DeRF with simple motion interpolation and progressive blur kernel estimation.
+With the emergence of the Gaussian splatting [42], researches such as [43, 44, 45] have also been published in this field, demonstrating that research remains active and ongoing.
+One of the most actively researched areas among those mentioned earlier is NeRF from blurry images, which often occurs when users take pictures with their own devices.
+However, as we mentioned in Section I, the experimental setup of using only 20$\sim$30 blurry images, as in previous studies, is not practical.
+If we assume a scenario where users only have access to blurry images, it is more realistic to consider that only a few images are available for a specific scene and all of those images are blurry.
+Therefore, we propose a more practical scenario by combining DeRF and the sparse view setting, thereby enhancing real-world applicability.
+II-C Radiance Fields from Sparse View
 There have been a lot of works to apply the neural representation in more pragmatic scenarios as the importance of VR and AR technologies increased, such as fast rendering, efficient sampling on rays, scene editing, denoising, and training from sparse view.
@@
 Sparse view images incur the inherent drawback of neural networks in that the network is more likely to be overfitted to the given data distribution.
-This leads to inconsistent scene geometry in the mapped representations, typically manifested as incorrectly predicted structural information, such as elongated density artifacts in the rendered color and depth images from novel views.
+This leads to lack of the information for dense scene geometry in the mapped representations, typically manifested as incorrectly predicted structure, such as elongated density artifacts in the rendered color and depth images from novel views.
 Several approaches have mitigated this issue involving additional prior knowledge or out-of-distribution data.
@@
 FreeNeRF [4] tries to alleviate the overfitting problem based on an optimization perspective, imposing some restrictions on the frequency level.
-In addition to sparse view settings, establishing NeRF from degraded images is recently emerging since the ideal training condition in images for NeRF often breaks in real-world scenarios.
-RawNeRF [43] denoises the internal noise of the camera sensor to construct high dynamic range (HDR) radiance fields from dark raw images and controls the camera exposure.
-Similarly, NaN [44] deals with burst noise in images, generating denoised images based on IBRNet [45], which is another image-based rendering approach.
-For more practical applications of NeRF in real-world, DeblurNeRF [1] firstly attempts to deal with two types of blur degradation in images, blur from camera motion and defocus, constructing deblurred neural radiance fields (DeRF) from only blurry images.
-They imitate the blurring process integrating the concept of blind deblurring in image deblurring with the NeRF system, modeling the blur kernel as pixel-wise independent ray transformation and composition weights to approximate the blurring process.
-Another representative approach is proposed by DP-NeRF[2], which imposes physical consistency across the images by modeling the blur kernel as the 3D rigid transformation of rays depending on each view, to approximate the actual blurring process in the camera more precisely.
-Recently several approaches [46, 47, 3] are also proposed in succession, attempting to improve the quality of the constructed DeRF.
-One of the most actively researched areas among those mentioned earlier is NeRF from blurry images, which often occurs when users take pictures with their own devices.
-However, as we mentioned in Section I, the experimental setup of using only 20$\sim$30 blurry images, as in previous studies, is not practical.
-If we assume a scenario where users only have access to blurry images, it is more realistic to consider that only a few images are available for a specific scene and all of those images are blurry.
-Therefore, we propose a more practical scenario by combining DeRF and the sparse view setting, thereby enhancing real-world applicability.
+There have been various studies attempting to address the challenge of accurate geometry prediction in sparse-view settings by utilizing depth priors obtained from pretrained depth estimation networks or point clouds generated by structure-from-motion (SfM).
+For example, SparseNeRF [52] tackled this issue by distilling depth ranking information, while SPARF [53] incorporated geometric warping to refine the depth priors for unseen views.
+DSNeRF [54] leveraged point clouds obtained from SfM as depth priors and Dense-Depth-Prior [55] further employed depth completion techniques to reduce the sparsity of the SfM point cloud, resulting in denser depth priors.
+StructNeRF [56] incorporate superpixel segmentation to refine the depth priors even more effectively.
+While these approaches attempt to address the lack of information by using depth priors, our proposed practical experimental settings that use the sparse blurry images reveal significant limitations.
+As mentioned earlier, the presence of blur in each image makes predicted depth inaccurate and the inconsistent across the multiple views, which hinders their usability as a prior.
 III preliminary
@@
 where $\hat{B}_{c}$, $\hat{B}_{f}$, and $B$ are expected coarse, fine, and ground truth blurred color of the ray r, respectively.
-The blur kernels are representatively modeled as a different type of transformation in each paper, [1, 2], which we will describe in the next paragraph.
-After the joint training of the approximated blur kernel and NeRF simultaneously, they can render the clean neural radiance fields, which we refer to as deblurred neural radiance fields (DeRF), by evaluating only NeRF thanks to the theoretical basis of blind deblurring.
+The blur kernels are representatively modeled as a different type of transformation in each paper, [1, 2].
 Figure 1: Simple illustration for different blur kernel modeling of (a) Deblur-NeRF [1] and (b) DP-NeRF [2].
 The main difference between the two kernels is the consistency between transformed rays induced from the blur kernel.
-Unlike Deblur-NeRF [1], which independently predict blurring rays for each pixel, DP-NeRF [2] predicts camera motion that makes the blur of each image and shares the motion across the entire pixels in the same image.
 III-B Blur Kernels in Deblurred Neural Radiance Fields
-The core difference of the blur kernels between the two papers is a consideration of 3D consistency across the entire pixels in each image as we describe in Figure 1.
+The core difference of the blur kernels between the two papers [1, 2] is a consideration of 3D consistency across the entire pixels in each image as we illustrate in Figure 1.
 Deblur-NeRF [1] designs the blur kernel as a deformable sparse kernel (DSK), which consists of the $n$ number of transformations depending on embedded latent features from each view and location of image pixels.
@@
 They utilize the physical intuition of actual blurred image acquisition in the camera process as an additional prior for the DeRF, to impose the constraints for constructing radiance fields while preserving 3D consistency.
+To ensure 3D consistency, unlike Deblur-NeRF [1], they made the blur process shared within a single image and defined it as a learnable set of camera motions.
 To directly model the actual camera motion as a 3D rigid transformation, they introduce scene-wise $SE(3)$ fields inspired by [29], [58] and Rodrigues’ formula [59].
@@
 In addition to modeling the blurring process with rigid blur kernel (RBK), [2] proposes an adaptive weight proposal network (AWP) based on the internal correlation between transformed rays and motion axis to predict the adaptive composition weights $\tilde{m}_{q}$ for each pixel, which complements the effect on the blur derived from the depth difference.
-In this paper, we present a novel approach that includes geometric constraints and a perceptual prior through extensive experiments based on these two types of blur kernels from Ddeblur-NeRF [1] and DP-NeRF [2], which can be regarded as flexible and rigid kernels, respectively.
+In this paper, we propose a novel approach that consists of geometric constraints and a perceptual prior through extensive experiments with these two types of blur kernels from Deblur-NeRF [1] and DP-NeRF [2].
 The conducted experiments demonstrate the effectiveness of Sparse-DeRF on both kernels.
-Furthermore, the results also reveal a trade-off for the flexibility of the kernels, depending on the specific properties inherent in the diverse scenes.
 IV Sparse-DeRF
 We find that it is not effective to construct the radiance fields based on naive NeRF, Deblur-NeRF [1], or DP-NeRF [2] from sparse view setting as we present in Section V-E.
-Although the DeRF usually recovers the high-frequency detail better than the naive NeRF in given views, it has become more fragmented in novel view synthesis, generating inaccurate scene geometry due to the complex joint optimization.
+Although the DeRF usually recovers the high-frequency detail better than the naive NeRF in given views, it has become more fragmented in novel view synthesis, generating inaccurate scene geometry due to the complex joint optimization and overfitting.
 The geometric error is represented as mapped RGB textures resembling painted walls in near or far depth and elongated density artifacts, which reveal challenges associated with accurate depth value prediction.
@@
 Figure 2 illustrates the overall architecture of the Sparse-DeRF in detail, where (a),(b), and (c) of the Figure 2 indicate each main component of the Sparse-DeRF, respectively.
-Geometric constraints consist of surface smoothness (SS) and modulated gradient scaling (MGS) as we describe in Section IV-A and Section IV-B.
-A perceptual prior consists of a perceptual distillation (PD) which is described in Section IV-C.
+Geometric constraints consist of surface smoothness (SS) and modulated gradient scaling (MGS), and a perceptual prior consists of a perceptual distillation (PD).
 Figure 2: Overall architecture of the Sparse-DeRF.
-Each color of integrated ray $\textbf{r}_{io}$ describes where the rays come from.
-$\hat{d}_{\Theta}(\textbf{r}_{i\sim i+k,j\sim j+k})$ indicates the rendered depth patch with $k\times k$ size.
-Note that, $k$ is set to $S_{ptc}$ as we described in the paper.
-The main component of the Sparse-DeRF consists of (a) surface smoothness (SS), (b) perceptual distillation (PD), and (c) modulated gradient scaling (MGS), which are indicated as gray, pink, and green colored components.
+The main component of Spare-DeRF consists of three components, Surface Smoothness (SS), Modulated Gradient Scaling (MGS), and Perceptual Distillation (PD), which are denoted as (a), (b), (c) in the figure.
+Note that, NeRF network is shared to predict the color of each sampled points along the ray in hidden rays, integrated unobserved rays and rays for patch rendering.
 Figure 3: Simple illustration of unobserved rays in our method, which consists of unseen rays and hidden rays.
@@
 6, generates the ray transformation to approximate the color composition process of the blur, regardless of the kernel types, as demonstrated in Deblur-NeRF [1] and DP-NeRF [2].
-Motivated by the commonality in the color composition process across various blur kernels, the kernel-induced transformed rays $\textbf{r}_{q}$ are introduced as additional unobserved rays, which are referred to as hidden rays, to enforce depth smoothness constraint.
-As the hidden rays $\textbf{r}_{q}$ are not directly presented in the training data for the DeRF system, they serve as supplementary out-of-distribution data for imposing depth smoothness, similar to the aforementioned unseen rays.
-In addition, the incorporation of depth smoothness on hidden rays effectively addresses geometric inconsistency across the 3D space of blurry images within the specified training view.
-This capability stems from the broader coverage of the estimated hidden rays, facilitated by the implicit inclusion of camera motion information in blurry images.
+Motivated by the commonality in the color composition process across various blur kernels, these kernel-induced transformed rays $\textbf{r}_{q}$ are introduced as additional unobserved rays, which are referred to as hidden rays, to enforce depth smoothness constraint to the blur kernel.
+As the hidden rays $\textbf{r}_{q}$ are not directly presented in the training data for the DeRF system, they serve as supplementary data for imposing depth smoothness regularization effect to the blur kernel and stabilize the complex optimization of blur kernel and radiance fields.
+In addition, the incorporation of depth smoothness on hidden rays effectively help alleviate the geometric inconsistency across the 3D space of blurry images within the specified training view.
 Therefore, our integrated unobserved rays for depth smoothness are defined as an integrated set of unseen rays and hidden rays as
@@
 Therefore, the gradient scaling seems to be more necessary to the radiance fields from the sparse view setting.
+In addition, we believe the proper gradient scaling can be an implicit geometric guidance in sparse view setting.
 Although we tried to apply the technique to the Sparse-DeRF, it does not properly works as demonstrated in the appendix.
@@
 [60] briefly mentioned the determinant of the jacobian as an additional scaling factor for the value of $J$, but they did not experiment on it.
-Moreover, even if the additional scaling factor were applied, the shape of the scaling function would remain unchanged and simply in the form of the square function, which does not allow for flexible adaptation to the arrangement of scene components.
-Therefore, we modulate the shape of the scaling function $J$ to adaptively reflect the scene arrangements and be suitable for NDC.
+Moreover, even if the additional scaling factor were applied, the shape of the scaling function would remain unchanged and simply in the form of the square function, which does not allow for flexible adaptation to the arrangement of scene components or guide the rough geometry of the radiance fields.
+Therefore, we modulate the shape of the scaling function $J$ to adaptively reflect the scene arrangements and guide the rough geometry while suitable for NDC, since we use NDC for our experiments.
 Our novel gradient scaling function $\hat{J}$ is designed based on three conditions.
 First, the function should increase from zero at the camera origin, which is a critical condition to avoid the local minima in the initial training phase we mentioned before.
-Second, the function should be not zero in the far distance, which is set to $1$ in our NDC environment, to ensure the NeRF training.
+Second, the function should not be zero in the far distance, which is set to $1$ in our NDC environment, to ensure the NeRF training.
+Second, the function should not be zero in the far distance, which is set to $1$, to ensure the NeRF training.
 Finally, the function is designed to increase and decrease only once within a given depth range, which makes the function not fluctuate, since it is an intuitively reasonable scenario considering the goal of MGS that alleviates incorrect density mapping in near distance.
@@
 This inconsistency occurs due to the ill-posed property of image deblurring and independent deblurring processing across multi-view images, which generates incoherent deblurred results.
-In the appendix, we additionally address this issue and present the qualitative comparison of pre-deblurred images and reference images, which are estimated from the DP-NeRF [2] trained with the full view, to reveal the geometric inconsistency issue.
-To overcome the intrinsic inconsistency of pre-deblurred images, we address the pre-filtered images as a perceptual prior, which transfers the deblurred texture information by extracting features from rendered patches and deblurred images with a pre-trained feature extractor.
+In Section V-F6, we additionally address this issue and present the qualitative comparison of pre-deblurred images and reference images, which are estimated from the DP-NeRF [2] trained with the full view, to reveal the geometric inconsistency issue.
+To overcome the intrinsic geometric inconsistency of pre-deblurred images, we address the pre-deblurred images as a perceptual prior, which transfers the only deblurred texture information by extracting features from rendered patches and deblurred images with a pre-trained feature extractor.
 Figure 5 simply illustrates the pipeline of our perceptual distillation module.
+To the best of our knowledge, utilizing pre-deblurred images as an additional prior in deblurring radiance fields is a first attempt.
 Specifically, we first generate deblurred images $\bar{I}$ for blurry training images $I$ by exploiting a pre-trained image deblurring network $\Theta_{D}$ to prepare the feature extraction as
@@
 where $\hat{\varphi}_{ptc}$ and $\bar{\varphi}_{ptc}$ indicate extracted features from each color patches, respectively.
-Note that, the color patch $\hat{C}_{ptc}$ is rendered by forwarding the NeRF MLPs without blur kernel to perceptually transfer the pre-filtered texture information to the implicit clean radiance fields, which is indicated as patch radiance in Figure 2.
+Note that, the color patch $\hat{C}_{ptc}$ is rendered by forwarding the NeRF MLPs without blur kernel to perceptually transfer the pre-deblurred texture information to the implicit clean radiance fields, which is indicated as patch radiance in Figure 2.
 Then we apply the perceptual loss $\mathcal{L}_{pd}$ [61] to distill the feature information as
@@
 Scenarios for the Sparse-DeRF are assumed to be three kinds of settings, which are the DeRF from 2-view, 4-view, and 6-view settings unlike existing common sparse-view NeRFs, which usually use 3-view, 6-view, and 9-view settings.
-The reason is that when using more than 8 images, the joint optimization problem occurs more frequently under 9-view settings as shown in Figure 6.
+The reason is that the joint optimization problem occurs more frequently under 9-view settings as shown in Figure 6.
 The figure presents the graph of experiments in the Decoration scene where we varied the number of input sparse views from 2-view to 10-view.
-We also attach the quantitative results in the appendix due to the page limitation.
+We also attach the quantitative results in the appendix.
 Experimental results reveal that our sparse-view experimental settings are valid since the RegNeRF [6] with DP-kernel shows poor performances under the 9 views.
@@
 V-E Evaluation
-TABLE I: Average results of novel view synthesis for both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings.
+TABLE I: Average results on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings.
 Each color shading represents the best, second best, and third best result for each experimental setting, respectively.
@@
 Naive NeRF [19] 15.11 .2999 .5578 14.38 .2635 .6004 20.02 .5327 .4000 18.98 .4860 .4481 21.65 .5985 .3638 20.63 .5513 .4014
-MPR [56] + NeRF 15.16 .3006 .5595 14.38 .2594 .6019 20.00 .5381 .3956 18.89 .4829 .4484 21.72 .5999 .3629 20.60 .5513 .4010
+NeRF + MPR [65] 15.16 .3006 .5595 14.38 .2594 .6019 20.00 .5381 .3956 18.89 .4829 .4484 21.72 .5999 .3629 20.60 .5513 .4010
 Deblur-NeRF [1] 15.14 .2884 .5330 14.41 .2506 .5921 19.99 .5199 .3499 18.89 .4761 .4151 23.12 .6798 .2386 21.36 .6003 .3163
@@
 RegNeRF [6] (No kernel) 14.60 .2849 .5869 15.49 .2997 .5888 18.39 .4600 .4704 18.44 .4326 .4852 19.69 .5249 .4165 19.65 .4790 .4583
+RegNeRF [6] (No kernel) + MPR [65] 14.99 .2980 .5690 15.84 .3040 .5818 18.54 .4678 .4593 18.66 .4425 .4804 19.89 .5339 .4071 19.64 .4753 .4597
 RegNeRF [6] (w/DP-kernel) 13.24 .2162 .6062 12.76 .1836 .6447 16.44 .3657 .4826 13.59 .2221 .5887 21.60 .6162 .3046 18.25 .4439 .4326
@@
 In the Sparse-DeRF results, DN-kernel and DP-kernel denote the blur kernels proposed by the Deblur-NeRF [1] and DP-NeRF [2], respectively.
-MPR$+$NeRF denotes the naive NeRF model trained solely on color supervision from deblurred images by MPRNet [56], utilizing the reconstruction loss $\mathcal{L}_{recon}$ of Eq.5.
-The results of MPR$+$NeRF reveal interesting observations and marginal improvements in radiance fields when only employing pre-deblurred images as direct color supervision for training.
+NeRF$+$MPR denotes the naive NeRF model trained solely on color supervision from deblurred images by MPRNet [65], utilizing the reconstruction loss $\mathcal{L}_{recon}$ of Eq.5, which is samely utilized in RegNeRF+MPR experiment.
+The results of the experiment with MPR [65] reveal interesting observations and marginal improvements in radiance fields when only employing pre-deblurred images as direct color supervision for training.
+This tendency even prominent when observing the results of applying MPR [65] to both NeRF (2 row in TABLE I) and RegNeRF (6 row in TABLE I).
+Regardless of the presence or absence of a regularizer, the improvement is marginal and the performance of perceptual quality(LPIPS score) even decreases when a larger number of images is used due to the inconsistency as we mentioned above.
+In the RegNeRF+MPR experiments, we observe a slight increase in PSNR in the 2-view setting.
+In this case, we expect the somewhat cleaner image from the pre-deblurred image appears to be the advantage, since there are significant lack of information in the experimental environment.
+Nonetheless, perceptual quality significantly deteriorates with poor LPIPS scores, which demonstrates the pre-deblurred images are not quite helpful to enhance the quality of the reconstructed radiance fields.
 This tendency emphasizes the 3D inconsistency across the pre-deblurred images, as discussed in Section IV-C.
@@
 We attach the experimental results of the RegNeRF [6] from 2-view, 4-view, and 6-view settings with and without the blur kernel to demonstrate the difficulty of the complex joint optimization problem as we mentioned.
-TABLE II and Figure 8 present quantitative and qualitative results from 2-view, 4-view, and 6-view, respectively.
-The blur kernel employed is the rigid blur kernel of DP-NeRF [2], which is described as DP-kernel in the table and figures.
+The results of the RegNeRF w/ and w/o DP-kernel in TABLE I (5, 6, 8 row) and Figure 8 present quantitative and qualitative evidence of complex optimization from sparse view setting and shows the outstanding performance of our model.
 To help the reader compare the plausible appearance and dense geometry with the proposed Sparse-DeRF, we attach the rendered color and depth images of our model.
@@
 The proposed Sparse-DeRF presents high-quality rendered images with dense geometry and detailed high-frequency texture as shown in Figure 8.
-TABLE II: Average results for the complex optimization issue in both synthetic and real scenes from 2-view, 4-view, and 6-view settings.
-Each color shading represents the best and second results for each experimental setting, respectively.
-[ 2-view ] [ 4-view ] [ 6-view ]
-[ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ]
-Blur Kernel  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
-RegNeRF [6] $\times$ 14.60 .2849 .5869 15.49 .2997 .5888 18.39 .4600 .4704 18.44 .4326 .4852 19.69 .5249 .4165 19.65 .4790 .4583
-RegNeRF [6] DP-kernel 13.24 .2162 .6062 12.76 .1836 .6447 16.44 .3657 .4826 13.59 .2221 .5887 21.60 .6162 .3046 18.25 .4439 .4326
-Sparse-DeRF (Ours) DP-kernel 15.35 .2904 .5242 15.57 .3114 .5467 21.05 .5776 .2975 20.05 .5178 .3736 24.27 .7255 .2044 22.32 .6283 .2907
 Figure 8: Qualitative results on Pool, Tanabata, and Heron scene for complex optimization issue from 2-view, 4-view, and 6-view settings.
 (i), (ii), and (ii) denote the rendered images from the RegNeRF [6] with no blur kernel, RegNeRF with DP-kernel, and Sparse-DeRF(Ours) with DP-kernel, respectively.
-TABLE III: Ablation experimental results from 2-view, 4-view, and 6-view according to the proposed geometric constraints and perceptual prior, denoted in the table as surface smoothness (SS), modulated gradient scaling (MGS), and perceptual distillation (PD), respectively.
-We present separate results for our model with both types of kernels, DP-kernel [2] and DN-kernel [1].
+TABLE II: Average results of the ablation study for the comparison of the proposed geometric constraints and perceptual prior, denoted in the table as surface smoothness (SS), modulated gradient scaling (MGS), and perceptual distillation (PD).
+The experiement is conducted on novel view synthesis for both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings.
 Each color shading represents the best, second best, and third best result for each experimental setting, respectively.
@@
 We demonstrate the effectiveness of the Sparse-DeRF’s each component through comprehensive ablation studies, presenting both quantitative and qualitative results.
-In TABLE III and Figure 9, we present independent quantitative and qualitative results from SS, MGS, and PD, where each component is individually applied to observe the influence of each method.
+In TABLE II and Figure 9, we present independent quantitative and qualitative results from SS, MGS, and PD, where each component is individually applied to observe the influence of each method.
 The results include several combinations of the components to reveal the complement effect of each proposed method.
@@
 If MGS is not applied, we can see that the geometry of the scene is not accurately captured or there are many density artifacts.
-On the other hand, PD seems to be not effective without geometric constraints in 2-view and 4-view settings as we can figure out in TABLE III and Figure 9.
+On the other hand, PD seems to be not effective without geometric constraints in 2-view and 4-view settings as we can figure out in TABLE II and Figure 9.
 The reason is that the NeRF has more difficulty in predicting the correct geometry with only pre-deblurred images due to the inherent 3D inconsistency, which makes perceptual prior not effective.
@@
 We can demonstrate that a certain level of accurate geometry should be achieved before applying perceptual distillation.
-TABLE IV: Average quantitative comparison results between naive gradient scaling of [51] and our MGS in both synthetic and real scenes from 2-view, 4-view, and 6-view settings.
-The two types of kernels we utilize in this paper are indicated as DP-kernel and DN-kernel, the kernels of DP-NeRF [2] and Deblur-NeRF [1], respectively.
+TABLE III: Average results between naive gradient scaling of [60] and our MGS.
+The experiements are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings.
 Color shading represents the better result.
@@
 Note that we attach the rendered color and depth images from DP-NeRF [2] trained with full view as reference images.
+TABLE IV: Average results of the comparison with the model that uses depth-prior.
+The experiments are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings.
+Each color shading represents the best, second best results.
+[ 2-view ] [ 4-view ] [ 6-view ]
+[ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ]
+ PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 18.12 .4729 .4884 18.01 .4683 .5245 21.08 .6105 .4042 20.69 .5909 .4450 21.77 .6369 .3903 21.28 .6189 .4250
+Sparse-DeRF (w/DN-kernel) - Ours 15.52 .2966 .5291 15.53 .3112 .5515 20.57 .5565 .3354 19.98 .5231 .3871 23.32 .6903 .2379 22.15 .6248 .3030
+Sparse-DeRF(w/DP-kernel) - Ours 15.35 .2904 .5242 15.57 .3114 .5467 21.05 .5776 .2975 20.05 .5178 .3736 24.27 .7255 .2044 22.32 .6283 .2907
+Figure 11: Qualitative results on Puppet, Ball, and Basket scene from 2-view, 4-view, and 6-view settings with the SparseNeRF [52] that use depth-prior.
+Even with inaccurate depth priors, it can be helpful in capturing rough geometry in extreme scenarios like a 2-view setting.
+However, when using 4 or 6 images, the multi-view inconsistency of the predicted depth prior leads to rather inaccurate geometry predictions, especially in thin or detailed structures.
+Note that, reference image is rendered from DP-NeRF [2] trained with full-view images for comparison.
+TABLE V: Average results of the comparison depending on usage of the pre-deblurred images on various models.
+The experiements are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings.
+Each color shading represents the best, second, and third result for each experimental setting, respectively.
+[ 2-view ] [ 4-view ] [ 6-view ]
+[ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ]
+ PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+Naive NeRF [19] 15.11 .2999 .5578 14.38 . 2635 .6004 20.02 .5327 .4000 18.98 .4860 .4481 21.65 .5985 .3638 20.63 .5513 .4014
+NeRF + MPR [65] 15.16 .3006 .5595 14.38 .2594 .6019 20.00 .5381 .3956 18.89 .4829 .4484 21.72 .5999 .3629 20.60 .5513 .4010
+RegNeRF [6] 14.60 .2849 .5869 15.49 .2997 .5888 18.39 .4600 .4704 18.44 .4326 .4852 19.69 .5249 .4165 19.65 .4790 .4583
+RegNeRF [6] + MPR [65] 14.99 .2980 .5690 15.84 .3040 .5818 18.54 .4678 .4593 18.66 .4425 .4804 19.89 .5339 .4071 19.64 .4753 .4597
+SparseNeRF [52] 18.12 .4729 .4884 18.01 .4683 .5245 21.08 .6105 .4042 20.69 .5909 .4450 21.77 .6369 .3903 21.28 .6189 .4250
+SparseNeRF [52] + MPR [65] 18.14 .4850 .4531 17.65 .4501 .5126 21.67 .6623 .3381 20.88 .6165 .4132 22.47 .6995 .3156 21.53 .6479 .3935
+Sparse-DeRF (w/DN-kernel) - Ours 15.52 .2966 .5291 15.53 .3112 .5515 20.57 .5565 .3354 19.98 .5231 .3871 23.32 .6903 .2379 22.15 .6248 .3030
+Sparse-DeRF(w/DP-kernel) - Ours 15.35 .2904 .5242 15.57 .3114 .5467 21.05 .5776 .2975 20.05 .5178 .3736 24.27 .7255 .2044 22.32 .6283 .2907
+Figure 12: Qualitative results on Heron scene from 4-view setting for the comparison of NeRF [19], RegNeRF [6] the SparseNeRF [52], and ours depending on the pre-deblurred images by using MPRNet [65].
+Note that, reference image is rendered from DP-NeRF [2] trained with full-view images for comparison.
 V-F4 Comparison to Naive Gradient Scaling
-We present experimental results of quantitative and qualitative comparison between our proposed MGS and naive gradient scaling of [51] from 2-view, 4-view, and 6-view settings in TABLE IV and Figure 10.
+We present experimental results of quantitative and qualitative comparison between our proposed MGS and naive gradient scaling of [60] from 2-view, 4-view, and 6-view settings in TABLE III and Figure 10.
 We compare the effectiveness of our MGS for two types of kernels we utilize in our paper, which are the kernels of Deblur-NeRF [1] and DP-NeRF [2].
@@
 Quantitative results for the entire scene are attached in the appendix.
-V-F5 Inconsistency of Pre-deblurred Images
-As mentioned in Section I and IV-C, image deblurring is conducted independently for each image and presents inconsistent geometry in various regions due to its ill-posed property.
+Moreover, since experiments on the DeblurNeRF [1] dataset were conducted in NDC by default, we adopted hyperparameters of MGS to be optimized for NDC.
+However, since MGS provides geometric guidance for the scene, it naturally works in other coordinate systems as well.
+To validate this, we also included experimental results in the appendix demonstrating its effectiveness in the world coordinate system.
+V-F5 Challenges in Utilizing Depth Priors
+We further address this inconsistency issue in relation to depth priors.
+The different blur in each image lead to accurate and multi-view inconsistent depth estimation, which is not proper to be used as additional prior for predict radiance fields with clean geometric structure.
+In TABLE IV and Figure 11, we present the quantitative and qualitative results of the proposed Sparse-DeRF and SparseNeRF [52], which is the representative depth-prior based model.
+The results show that as the number of images increases, the improvement diminishes and scene geometry is distorted, revealing the issue of inconsistent depth priors.
+In extreme cases like a 2-view setting, where it is challenging to obtain even rough geometry information, even inaccurate depth maps can be somewhat helpful.
+In addition, since only two images are used, the inconsistency issue is less pronounced, resulting in relatively better structured geometry.
+However, in 4- or 6-view settings, relying on inconsistent depth prior can make the prediction of a clean and accurate scene geometry even more difficult, particularly for thin or detailed structures as we can observe in the emphasized region in Figure 11.
+In particular, the qualitative experimental results with 4 views further support the inconsistency issue, as the significant difference in perceptual quality, reflected by LPIPS, provides additional evidence.
+It is once again clear that qualitative evaluation and LPIPS performance are crucial in the deblurring task, while PSNR and SSIM are important, as mentioned in DP-NeRF [2].
+Furthermore, in the appendix, we included the experimental results of SparseNeRF [52] with depth maps predicted by Depth-Anything V2 [68], one of the state-of-the-art monocular depth estimation models, which also exhibits the experimental tendency and multi-view inconsistency as well.
+Therefore, to design a general sparse view model that works well across 2-, 4-, and 6-views, we adopted an experimental setup that does not utilize depth priors.
+Quantitative results for the entire scene are attached in the appendix.
+V-F6 Inconsistency of Pre-deblurred Images
+Image deblurring is conducted independently for each image and presents inconsistent geometry in various regions due to its ill-posed property.
 In Figure 13, we present the qualitative comparison to reveal the inconsistency of pre-deblurred images across the multi-view training images.
@@
 Comparing emphasized regions from each image, the pre-deblurred image shows relatively well-restored textures in each image, but the geometry is inconsistently restored and distorted across the multiple views.
-Such inconsistency adversely affects the learning of radiance fields since it is trained by pixel-wise color reconstruction loss.
-Furthermore, the performance of image deblurring varies even within a single image depending on the structural complexity of the local region, making it challenging to learn blur kernels.
-Due to these reasons, directly utilizing pre-filtered images for training radiance fields is difficult.
+Such inconsistency adversely affects the learning of radiance fields, which is crucial for perceptual quality, since it is trained by pixel-wise color reconstruction loss.
+In TABLE V, we present the comparison results of the NeRF [19], RegNeRF [6], and SparseNeRF [52] with and without the pre-deblurred images.
+The results from 4-view and 6-view also demonstrate that, it is not enough to reconstruct the clean radiance fields with high perceptual quality only depending on the pre-deblurring and depth estimation although they provide the abundant additional priors.
+To qualitatively demonstrate the observation, we included Figure 12, which compares the results trained on 4 view setting.
+While SparseNeRF [52] + MPR [65], which utilizes both pre-deblurred images and depth prior, achieved higher PSNR and SSIM scores numerically, the rendered results clearly show that our approach produces much sharper and more precise renderings.
+Due to these reasons, directly utilizing pre-deblurred images for training radiance fields is difficult.
 Therefore, perceptual distillation is introduced to transfer only the perceptual texture of pre-deblurred images to the radiance fields.
@@
 VI Limitations and Discussions
-Despite the remarkable enhancement in terms of 3D geometry and appearance, there are still several limitations.
+Despite the remarkable enhancement in terms of 3D geometry and appearance, there are several points for improvement.
 The first one is derived from the blur kernel itself, especially the relationship between the type of the blur kernel and the properties of each scene.
@@
 The second one is that we have to set the proper hyper-parameter for MGS to find the most effective function shape although MGS greatly improves the 3D geometry in the DeRF from sparse view.
-However, it is difficult to find an ideal function shape according to the arrangement of the object in the scene, especially as we mentioned above.
+However, it is difficult to find an ideal function shape according to the arrangement of the object in the scene.
 We handle these cases by setting the magnitude $\rho$ as a high value to only ignore the gradient in a very near distance region, which is attached to the appendix as detailed hyper-parameters per each scene.
@@
 The innate challenge of the 2-view setting is that sparse overlapped 3D space usually leads to inaccurate geometry, which is more likely to be mapped to be painted texture on the wall at the near or far depth regions.
-There is still room to improve the visual quality of the Sparse-DeRF and solve these problems through state-of-the-art generative methods such as diffusion models, which can be a great future direction for constructing the DeRF from sparse view.
+There is still room to improve the visual quality of the Sparse-DeRF and solve these problems through state-of-the-art generative methods similar to the work like Diffusionerf [69], which can be a great future direction for constructing the DeRF from sparse view.
+On the other hand, designing models that account for inconsistency to utilize additional data-deriven priors, such as depth and normal priors from pretrained models, could be an interesting direction for future work.
+Furthermore, leveraging real-time rendering models, like the recently emerging 3D Gaussian Splatting [42], could enhance practical applicability.
 VII Conclusion
@@
 As we mentioned in Section V-C in the main paper, we utilize the fixed unseen rays for training our model to fairly evaluate the performances across the extensive experiments since the randomness of unseen ray generation in the RegNeRF makes it hard to understand the effectiveness of each component.
-TABLE V: Image indices of each scene for training Sparse-DeRF.
+TABLE VI: Average results of the comparison with the model that uses depth-prior.
+The experiments are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings.
+Each color shading represents the best, second best, and third best results, respectively.
+[ 2-view ] [ 4-view ] [ 6-view ]
+[ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ]
+ PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 18.12 .4729 .4884 18.01 .4683 .5245 21.07 .6105 .4042 20.69 .5909 .4450 21.77 .6369 .3903 21.28 .6189 .4250
+SparseNeRF [52] + Depth-Any V2 [68] 18.24 .4909 .4479 16.35 .4262 .4768 21.78 .6679 .3354 18.87 .5406 .4016 22.56 .7034 .3141 19.44 .5659 .3845
+Sparse-DeRF (w/DN-kernel) - Ours 15.52 .2966 .5291 15.53 .3112 .5515 20.57 .5565 .3354 19.98 .5231 .3871 23.32 .6903 .2379 22.15 .6248 .3030
+Sparse-DeRF(w/DP-kernel) - Ours 15.35 .2904 .5242 15.57 .3114 .5467 21.05 .5776 .2975 20.05 .5178 .3736 24.27 .7255 .2044 22.32 .6283 .2907
+TABLE VII: Quantitative results for the entire scenes of real scenes from 2-view settings.
+Each color shading represents the best, second best, and third best result, respectively.
+Depth-Any V2 denotes the Depth-Anything V2 [68].
+Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
+PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 15.80 .3622 .5988 22.06 .7001 .3419 23.90 .6063 .3824 14.53 .3454 .5651 14.32 .3499 .5651 18.12 .4728 .4883
+SparseNeRF [52] + Depth-Any V2 [68] 15.93 .3923 .5060 23.02 .7419 .2972 23.64 .5968 .3957 13.94 .3086 .5469 14.71 .4150 .4938 18.24 .4909 .4479
+Sparse-DeRF (w/DN-kernel) - Ours 14.27 .2200 .6564 20.57 .5675 .3589 19.81 .3330 .4058 11.67 .1827 .5947 11.28 .1800 .6298 15.52 .2966 .5291
+Sparse-DeRF (w/DP-kernel) - Ours 14.10 .2116 .6413 18.97 .5206 .3417 20.32 .3488 .4056 12.00 .1943 .5902 11.36 .1769 .6422 15.35 .2904 .5242
+Real Scene Ball Basket Buick Coffee Decoration
+PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 21.62 .5588 .5092 16.32 .4636 .5488 18.32 .5356 .4754 24.82 .8225 .2958 14.85 .3325 .6020
+SparseNeRF [52] + Depth-Any V2 [68] 21.58 .5573 .5107 16.44 .4698 .5462 18.28 .5352 .4752 24.65 .8195 .3003 15.19 .3432 .5941
+Sparse-DeRF (w/DN-kernel) - Ours 20.07 .4612 .5290 13.62 .2678 .5835 14.38 .3182 .4913 19.51 .6048 .4146 13.09 .2173 .6256
+Sparse-DeRF (w/DP-kernel) - Ours 20.09 .4583 .5244 14.05 .2890 .5555 13.78 .2864 .5036 19.67 .6084 .4096 13.11 .2192 .6192
+Real Scene Girl Heron Parterre Puppet Stair Average
+PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 14.16 .4380 .5641 16.81 .3257 .5566 17.58 .3601 .5928 16.89 .4677 .5261 18.75 .3784 .5738 18.01 .4683 .5245
+SparseNeRF [52] + Depth-Any V2 [68] 14.65 .4600 .5513 16.90 .3258 .5597 17.25 .3481 .5983 16.66 .4684 .5234 18.33 .3614 .5861 16.35 .4262 .4768
+Sparse-DeRF (w/DN-kernel) - Ours 12.85 .3418 .5585 14.30 .1706 .5197 16.61 .2903 .5912 14.91 .3141 .5886 15.97 .1255 .6125 15.53 .3112 .5515
+Sparse-DeRF (w/DP-kernel) - Ours 13.11 .3707 .5486 14.17 .1669 .5285 16.19 .2527 .5908 15.25 .3277 .5739 16.30 .1343 .6133 15.57 .3114 .5467
+TABLE VIII: Quantitative results for the entire scenes of real scenes from 4-view settings.
+Each color shading represents the best, second best, and third best result, respectively.
+Depth-Any V2 denotes the Depth-Anything V2 [68].
+Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
+PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 17.15 .4271 .5516 23.50 .7531 .2994 26.63 .7091 .3114 18.93 .5838 .4389 19.14 .5865 .4195 21.07 .6105 .4042
+SparseNeRF [52] + Depth-Any V2 [68] 18.93 .5813 .3959 24.86 .8044 .2335 26.74 .7047 .2988 18.90 .6111 .3955 19.46 .6379 .3482 21.78 .6679 .3354
+Sparse-DeRF (w/DN-kernel) - Ours 16.91 .3693 .5357 25.79 .7872 .1801 25.48 .6042 .2405 16.58 .4705 .3793 18.10 .5514 .3414 20.57 .5565 .3354
+Sparse-DeRF (w/DP-kernel) - Ours 18.99 .4770 .4034 26.51 .8051 .1627 23.33 .4888 .2733 17.51 .5371 .3356 18.92 .5799 .3126 21.05 .5776 .2975
+Real Scene Ball Basket Buick Coffee Decoration
+PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 21.56 .5539 .5020 20.45 .5986 .4620 21.04 .6392 .4008 26.55 .8499 .2739 18.94 .5128 .5042
+SparseNeRF [52] + Depth-Any V2 [68] 21.74 .5587 .4973 20.55 .6007 .4611 21.17 .6474 .3987 26.47 .8504 .2696 18.86 .5097 .5031
+Sparse-DeRF (w/DN-kernel) - Ours 22.28 .5506 .4334 18.69 .5058 .3774 19.28 .5414 .3371 26.76 .8126 .2419 17.12 .4192 .4761
+Sparse-DeRF (w/DP-kernel) - Ours 23.39 .6010 .3806 20.41 .5451 .3305 19.48 .5531 .3251 27.77 .8364 .2049 16.33 .3914 .4950
+Real Scene Girl Heron Parterre Puppet Stair Average
+PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 17.96 .6579 .4445 19.08 .4873 .4793 20.91 .5239 .4808 18.56 .5409 .4444 21.83 .5446 .4574 20.69 .5909 .4450
+SparseNeRF [52] + Depth-Any V2 [68] 18.10 .6633 .4371 19.03 .4856 .4807 21.07 .5276 .4777 18.67 .5506 .4378 21.99 .5530 .4545 18.87 .5406 .4016
+Sparse-DeRF (w/DN-kernel) - Ours 16.95 .5975 .3823 19.04 .4558 .3315 20.74 .4916 .4323 18.11 .4616 .4202 20.80 .3948 .4386 19.98 .5231 .3871
+Sparse-DeRF (w/DP-kernel) - Ours 16.47 .5855 .3835 19.09 .4624 .3194 18.36 .3375 .4715 17.97 .4612 .4206 21.25 .4045 .4044 20.05 .5178 .3736
+TABLE IX: Quantitative results for the entire scenes of real scenes from 6-view settings.
+Each color shading represents the best, second best, and third best result, respectively.
+Depth-Any V2 denotes the Depth-Anything V2 [68].
+Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
+PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 17.17 .4286 .5402 23.86 .7601 .3021 28.04 .7612 .2772 19.99 .6216 .4212 19.78 .6131 .4106 21.77 .6369 .3903
+SparseNeRF [52] + Depth-Any V2 [68] 19.42 .6130 .3777 25.10 .8130 .2282 28.09 .7628 .2643 19.78 .6529 .3686 20.45 .6751 .3318 22.56 .7034 .3141
+Sparse-DeRF (w/DN-kernel) - Ours 18.33 .5144 .4004 26.63 .8184 .1513 28.16 .7344 .1755 22.06 .6988 .2245 21.43 .6855 .2380 23.32 .6903 .2379
+Sparse-DeRF (w/DP-kernel) - Ours 21.29 .6179 .3261 27.34 .8340 .1298 28.19 .7365 .1606 22.30 .7233 .2017 22.21 .7159 .2036 24.27 .7255 .2044
+Real Scene Ball Basket Buick Coffee Decoration
+PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 22.17 .5654 .4939 21.86 .6733 .4081 21.63 .6719 .3747 26.17 .8382 .2874 19.94 .5696 .4646
+SparseNeRF [52] + Depth-Any V2 [68] 22.14 .5661 .4896 21.96 .6746 .4099 21.66 .6754 .3736 26.21 .8373 .2847 20.04 .5763 .4615
+Sparse-DeRF (w/DN-kernel) - Ours 23.76 .6184 .3607 23.39 .7155 .2329 21.66 .6405 .2585 27.84 .8388 .2068 19.98 .5553 .3644
+Sparse-DeRF (w/DP-kernel) - Ours 24.80 .6643 .3200 23.32 .6789 .2455 21.27 .6234 .2662 28.91 .8525 .1902 19.69 .5440 .3672
+Real Scene Girl Heron Parterre Puppet Stair Average
+PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+SparseNeRF [52] 18.92 .7073 .4091 19.09 .4705 .4895 21.54 .5630 .4536 19.65 .5951 .4017 21.79 .5345 .4673 21.28 .6189 .4250
+SparseNeRF [52] + Depth-Any V2 [68] 19.40 .7156 .4050 19.17 .4731 .4900 21.68 .5701 .4500 19.80 .5995 .4000 21.84 .5369 .4657 19.44 .5659 .3845
+Sparse-DeRF (w/DN-kernel) - Ours 18.86 .6860 .2769 19.63 .5011 .2992 22.15 .5703 .3787 20.81 .5882 .3149 23.42 .5340 .3374 22.15 .6248 .3030
+Sparse-DeRF (w/DP-kernel) - Ours 18.96 .6887 .2744 19.59 .5051 .2808 22.57 .5994 .3317 20.60 .5756 .3124 23.48 .5509 .3189 22.32 .6283 .2907
+Figure 15: Additional qualitative ablation results on Puppet, Ball, and Basket scene from 2-view, 4-view, and 6-view settings, which are the same scene in the main paper.
+The results show the comparison between SparseNeRF [52] and SparseNeRF [52] trained with the depth maps predicted from Depth-Anything V2 [68], which are indicated as reference number [52] and Depth-Any V2, respectively.
+Appendix B Additional Analysis and Experimental Results
+B-A Models with Depth Prior
+We present the quantitative results of our model and SparseNeRF [52] in TABLE VIVIIVIII, and IX.
+Additionally, we include the results of SparseNeRF [52] trained with the more recent state-of-the-art monocular depth estimation model, Depth-Anything V2 [68], in these tables and present the qualitative comparison in Figure 15.
+Interestingly, the results show that SparseNeRF [52] using depth maps predicted by Depth-Anything V2 [68] yields lower PNSR and SSIM scores but a higher LPIPS score in real-world scenes compared to the originally adopted depth maps from DPT [70].
+We believe that this degradation occurs because Depth-Anything V2 [68] predicts more accurate depth maps even from blurred images in real-world data.
+However, this improvement in depth estimation unexpectedly exacerbates multi-view inconsistency among the predicted depth maps, which in turn negatively affects the optimization process of the radiance field in real scenes.
+As a consequence, the multi-view inconsistency issue, which becomes particularly noticeable across four or more views, remains unresolved.
+This degradation is also evident in the lower quality of the rendered images in Figure 15.
+TABLE X: Average results of MGS experiments on world coordinate systems.
+The experiments are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings.
+Color shading represents the best and second results, respectively.
+[ 2-view ] [ 4-view ] [ 6-view ]
+[ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ] [ Synthetic Scene ] [ Real Scene ]
+Blur Kernel Gradient Scaling  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 14.46 .2614 .6258 13.26 .1797 .6732 18.46 .4554 .4691 16.92 .3679 .4570 19.04 .4822 .4408 19.38 .4760 .3740
+DP-kernel + Naive [60] 14.41 .2579 .6204 13.20 .1789 .6680 18.40 .4506 .4706 17.05 .3654 .4544 19.01 .4830 .4425 19.12 .4575 .3831
+DP-kernel + MGS 14.49 .2564 .5974 15.22 .2880 .5362 18.54 .4493 .4300 17.91 .4061 .4019 19.10 .4781 .3963 19.79 .4923 .3470
+Figure 16: Additional qualitative ablation results on Basket, Coffee, and Decoration scene from 2-view, 4-view, and 6-view settings, which are trained on world coordinate system.
+The results demonstrate that our proposed MGS guides the geometry of the scenes and works properly even in the world coordinate system.
+B-B Geometric Guidance of Modulated Gradient Scaling
+We initially explored a gradient scaling function that could operate in normalized device coordinate (NDC).
+However, inspired by the idea that gradient scaling could serve as geometric guidance for the scene, we proposed MGS.
+To demonstrate that MGS can function as a structural clue regardless of the coordinate systems, we present TABLE X along with Figure 16.
+The experiments were conducted in the world coordinate system, using the same MGS parameters for each scene as presented in TABLE XI to match the scaling function shape in NDC for fair comparison.
+We converted the distances of ray samples along the rays in the world coordinate system to the NDC depth scale and used them as $\delta_{\textbf{s}_{i}}$ in Eq.21 of the main paper.
+Note that, other than this scaling adjustment, all settings for training the radiance fields remained in the world coordinate system.
+The kernel employed for the experiments is DP-kernel.
+The results show that our proposed MGS achieves superior performance even in the world coordinate system.
+These findings validate that MGS is not limited to a specific coordinate system but effectively captures the arrangement of scene components, facilitating the learning of radiance fields as intended.
+We additionally attach the quantitative results for entire scenes in Section D-D.
+TABLE XI: Image indices of each scene for training Sparse-DeRF.
 Real Scene Ball Basket Buick Coffee Decoration Girl Heron Parterre Puppet Stair
@@
 $6$-view 2, 14, 17, 21, 23, 29 1, 3, 14, 19, 28, 33 1, 5, 10, 15, 20, 23 1, 7, 11, 18, 22, 27 7, 13, 20, 23, 27, 31
-TABLE VI: Hyper parameters of modulated gradient scaling (MGS) for entire scenes of the Deblur-NeRF [1] dataset. $\rho$ and $\eta$ denote the magnitude and period of sine function, respectively.
+TABLE XII: Hyper parameters of modulated gradient scaling (MGS) for entire scenes of the Deblur-NeRF [1] dataset. $\rho$ and $\eta$ denote the magnitude and period of sine function, respectively.
 Synthetic Scene Cozyroom Factory Pool Tanabata Trolley
@@
 $\eta$ 1.2 0.67 1.75 0.67 0.5 0.5 0.5 0.5 1.75 0.5
-Appendix B Additional Implementation Details
-B-A Training Scene Indices
-For fair comparison in future research, we present the image indices of each scene for training the Sparse-DeRF in TABLE V.
+Appendix C Additional Implementation Details
+C-A Training Scene Indices
+For fair comparison in future research, we present the image indices of each scene for training the Sparse-DeRF in TABLE XI.
 The indices are manually selected from the training images of each scene as we mentioned in the main paper.
 Note that, the indices of the 2-view and 4-view settings are subsets of the 6-view setting.
-B-B Parameters for Entire Scenes
-In TABLE VI, we present the hyper-parameters of MGS for entire scenes, which consist of period $\eta$ and magnitude $\rho$ of the sine function.
+C-B Parameters for Entire Scenes
+In TABLE XII, we present the hyper-parameters of Modulated Gradient Scaling (MGS) for entire scenes, which consist of period $\eta$ and magnitude $\rho$ of the sine function.
 As we indicate in Section IV-B in the main paper, we set the magnitude $\rho$ as high value to only ignore the gradient in very near distance regions for scenes such as Buick, Puppet, Cozyroom, Factory, Pool, and Trolley.
 Please refer to the figure of the function shape depending on the parameters in the main paper.
-Appendix C Additional Quantitative Results
-C-A Quantitative Results for Entire Scenes
-We present the comprehensive experimental results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE VIIVIII, and IX, respectively.
-The results of the Sparse-DeRF in the TABLE VIIVIII, and IX are representative results as we presented in Section V-E1 in the main paper.
-In addition, we present more detailed ablation results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XXI, and XII, respectively.
+Appendix D Additional Quantitative Results
+D-A Quantitative Results for Entire Scenes
+We present the comprehensive experimental results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XIIIXIV, and XV, respectively.
+In the tables, DP-kernel and DN-kernel denote the kernel of [2] and [1], respectively.
+The results of the Sparse-DeRF in the TABLE XIIIXIV, and XV are representative results as we presented in Section V-E1 in the main paper.
+In addition, we present more detailed ablation results for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XVIXVII, and XVIII, respectively.
 As we mentioned before, we also present the separated ablation results according to the type of the kernel from DP-NeRF [2] and Deblur-NeRF [1].
-C-B Complex Optimization Problem for Entire Scenes
-We present the comprehensive experimental results of the complex optimization problem for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XIVXV, and XVI, respectively.
-In addition, we also present the quantitative results of the experiments on the Decoration scene that varies the number of sparse views in TABLE XIII as we mentioned in the main paper.
-C-C Comparison to Naive Gradient Scaling for Entire Scenes
-We present the quantitative comparison results for entire scenes that compare our modulated gradient scaling (MGS) and naive gradient scaling of [51] from 2-view, 4-view, and 6-view settings in TABLEXVIIXVIII, and XIX, respectively.
-TABLE VII: Quantitative results of novel view synthesis for the entire scene of synthetic and real scenes obtained from 2-view settings.
-Each color shading represents the best, second best and third best result, respectively.
-DP-kernel and DN-kernel denote the kernel of [2] and [1].
+D-B Complex Optimization Problem for Entire Scenes
+We present the comprehensive experimental results of the complex optimization problem for entire scenes from 2-view, 4-view, and 6-view settings in TABLE XXXXI, and XXII, respectively.
+In addition, we also present the quantitative results of the experiments on the Decoration scene that varies the number of sparse views in TABLE XIX as we mentioned in the main paper.
+D-C Comparison to Naive Gradient Scaling for Entire Scenes
+We present the quantitative comparison results for entire scenes that compare our modulated gradient scaling (MGS) and naive gradient scaling of [60] from 2-view, 4-view, and 6-view settings in TABLEXXIIIXXIV, and XXV, respectively.
+D-D World Coordinate System Experiment for Entire Scenes
+We present the quantitative comparison results for entire scenes that include MGS experiments on the world coordinate system from 2-view, 4-view, and 6-view settings in TABLEXXVIXXVII, and XXVIII, respectively.
+The results demonstrate the ability of our modulated gradient scaling (MGS) as geometric guidance.
+TABLE XIII: Quantitative results for the entire scene of synthetic and real scenes from 2-view settings. Each color shading represents the best, second best and third best result, respectively.
 Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
@@
 Sparse-DeRF (w/DP-kernel) - Ours 13.11 .3707 .5486 14.17 .1669 .5285 16.19 .2527 .5908 15.25 .3277 .5739 16.30 .1343 .6133 15.57 .3114 .5467
-TABLE VIII: Quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 4-view settings.
-Each color shading represents the best, second best and third best result, respectively.
-DP-kernel and DN-kernel denote the kernel of [2] and [1].
+TABLE XIV: Quantitative results for the entire scenes of synthetic and real scenes from 4-view settings. Each color shading represents the best, second best and third best result, respectively.
 Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
@@
 PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
-Naive NeRF [19] 21.61 .5043 .4992 17.23 fv.4384 .4746 19.20 .5508 .3579 23.21 .7341 .3281 14.78 .3299 .5640
+Naive NeRF [19] 21.61 .5043 .4992 17.23 .4384 .4746 19.20 .5508 .3579 23.21 .7341 .3281 14.78 .3299 .5640
 MPR [65]+NeRF 21.46 .5046 .4992 16.98 .4260 .4766 19.39 .5486 .3629 22.98 .7267 .3284 14.88 .3460 .5532
@@
 Sparse-DeRF (w/DP-kernel) - Ours 16.47 .5855 .3835 19.09 .4624 .3194 18.36 .3375 .4715 17.97 .4612 .4206 21.25 .4045 .4044 20.05 .5178 .3736
-TABLE IX: Quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 6-view settings.
+TABLE XV: Quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 6-view settings.
 Each color shading represents the best, second best and third best result, respectively.
@@
 Sparse-DeRF (w/DP-kernel) - Ours 18.96 .6887 .2744 19.59 .5051 .2808 22.57 .5994 .3317 20.60 .5756 .3124 23.48 .5509 .3189 22.32 .6283 .2907
-TABLE X: Ablation quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 2-view settings.
+TABLE XVI: Ablation quantitative results for the entire scenes of synthetic and real scenes from 2-view settings.
 Each color shading represents the best, second best and third best result, respectively.
-DP-kernel and DN-kernel denote the kernel of [2] and [1].
 Kernel Type Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
@@
 12.85 .3418 .5585 14.30 .1706 .5197 16.61 .2903 .5912 14.91 .3141 .5886 15.97 .1255 .6125 15.53 .3112 .5515
-TABLE XI: Ablation quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 4-view settings.
+TABLE XVII: Ablation quantitative results for the entire scenes of synthetic and real scenes obtained from 4-view settings.
 Each color shading represents the best, second best and third best result, respectively.
-DP-kernel and DN-kernel denote the kernel of [2] and [1].
 Kernel Type Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
@@
 16.95 .5975 .3823 19.04 .4558 .3315 20.74 .4916 .4323 18.11 .4616 .4202 20.80 .3948 .4386 19.98 .5231 .3871
-TABLE XII: Ablation quantitative results of novel view synthesis for the entire scenes of synthetic and real scenes obtained from 6-view settings.
+TABLE XVIII: Ablation quantitative results for the entire scenes of synthetic and real scenes obtained from 6-view settings.
 Each color shading represents the best, second best and third best result, respectively.
-DP-kernel and DN-kernel denote the kernel of [2] and [1].
 Kernel Type Synthetic Scene Factory Cozyroom Pool Tanabata Trolley Average
@@
 18.86 .6860 .2769 19.63 .5011 .2992 22.15 .5703 .3787 20.81 .5882 .3149 23.42 .5340 .3374 22.15 .6248 .3030
-TABLE XIII: Quantitative results of the RegNeRF [6] with and without the blur kernel from 2-view to 10-view settings.
+TABLE XIX: Quantitative results of the RegNeRF [6] with and without the blur kernel from 2-view to 10-view settings.
 For comparison, the blur kernel of the DP-NeRF [2] is employed, which is denoted as DP-kernel.
@@
 RegNeRF [6] DP-kernel .7520 .7191 .7089 .6992 .6932 .6547 .5277 .4178 .2804
-TABLE XIV: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 2-view settings.
+TABLE XX: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 2-view settings.
 Each color shading represents the best and second results for each experimental setting, respectively.
@@
 Sparse-DeRF(Ours) DP-kernel 13.11 .3707 .5486 14.17 .1669 .5285 16.19 .2527 .5908 15.25 .3277 .5739 16.30 .1343 .6133 15.57 .3114 .5467
-TABLE XV: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 4-view settings.
+TABLE XXI: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 4-view settings.
 Each color shading represents the best and second results for each experimental setting, respectively.
@@
 Sparse-DeRF(Ours) DP-kernel 16.47 .5855 .3835 19.09 .4624 .3194 18.36 .3375 .4715 17.97 .4612 .4206 21.25 .4045 .4044 20.05 .5178 .3736
-TABLE XVI: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 6-view settings.
+TABLE XXII: Entire experimental results for the complex optimization issue in both synthetic and real scenes from 6-view settings.
 Each color shading represents the best and second results for each experimental setting, respectively.
@@
 Sparse-DeRF(Ours) DP-kernel 18.96 .6887 .2744 19.59 .5051 .2808 22.57 .5994 .3317 20.60 .5756 .3124 23.48 .5509 .3189 22.32 .6283 .2907
-TABLE XVII: Quantitative comparison between naive gradient scaling of [51] and our MGS for entire scenes of synthetic and real scenes obtained from 2-view settings.
-The experimental results are presented according to the two types of kernel we utilize in this paper, which are kernels of the DP-NeRF [2] and Deblur-NeRF [1].
-Each kernel is denoted as DP-kernel and DN-kernel, respectively.
+TABLE XXIII: Quantitative comparison between naive gradient scaling of [60] and our MGS for entire scenes of synthetic and real scenes from 2-view settings.
+The experimental results are presented according to the two types of kernel we utilize in this paper.
 Color shading represents the better result.
@@
 DP-kernel $+$ MGS 12.89 .3465 .5645 14.48 .1677 .5209 16.20 .2511 .5927 14.48 .2941 .5819 15.81 .1167 .6039 15.46 .3035 .5465
-TABLE XVIII: Quantitative comparison between naive gradient scaling of [51] and our MGS for entire scenes of synthetic and real scenes obtained from 4-view settings.
-The experimental results are presented according to the two types of kernel we utilize in this paper, which are the kernels of DP-NeRF [2] and Deblur-NeRF [1].
-Each kernel is denoted as DP-kernel and DN-kernel, respectively.
+TABLE XXIV: Quantitative comparison between naive gradient scaling of [60] and our MGS for entire scenes of synthetic and real scenes from 4-view settings.
+The experimental results are presented according to the two types of kernel we utilize in this paper.
 Color shading represents the better result.
@@
 DP-kernel $+$ MGS 17.03 .6045 .3636 18.90 .4526 .3185 18.38 .3323 .4575 17.89 .4408 .4333 20.33 .3416 .4297 19.75 .4980 .3737
-TABLE XIX: Quantitative comparison between naive gradient scaling of [51] and our MGS for entire scenes of synthetic and real scenes obtained from 6-view settings.
-The experimental results are presented according to the two types of kernel we utilize in this paper, which are the kernels of DP-NeRF [2] and Deblur-NeRF [1].
-Each kernel is denoted as DP-kernel and DN-kernel, respectively.
+TABLE XXV: Quantitative comparison between naive gradient scaling of [60] and our MGS for entire scenes of synthetic and real scenes from 6-view settings.
+The experimental results are presented according to the two types of kernel we utilize in this paper.
 Color shading represents the better result.
@@
 DP-kernel $+$ MGS 18.77 .6833 .2703 19.58 .4998 .2856 22.27 .5801 .3440 20.48 .5790 .3043 23.23 .5468 .3149 22.07 .6212 .2889
+TABLE XXVI: Quantitative comparison results for entire scenes from a 2-view setting that include MGS experiments on the world coordinate system. Color shading represents the best and second result.
+Synthetic Scene Gradient Scaling Factory Cozyroom Pool Tanabata Trolley Average
+Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 15.17 .2710 .6140 13.48 .2441 .6758 15.24 .1592 .8082 15.00 .3469 .4822 13.43 .2859 .5489 14.46 .2614 .6258
+DP-kernel $+$ Naive [60] 15.10 .2599 .6074 13.70 .2598 .6662 15.41 .1729 .7816 14.85 .3403 .4823 13.00 .2566 .5643 14.41 .2579 .6204
+DP-kernel $+$ MGS 15.30 .2682 .6026 13.12 .2199 .6755 15.82 .1592 .6917 14.97 .3626 .4710 13.25 .2723 .5460 14.49 .2564 .5974
+Real Scene Gradient Scaling Ball Basket Buick Coffee Decoration
+Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 19.25 .4315 .5415 13.59 .2255 .5975 10.86 .0918 .6892 16.06 .4164 .6220 11.06 .0899 .7539
+DP-kernel $+$ Naive [60] 19.38 .4441 .5298 13.23 .2343 .5941 11.02 .1003 .6903 15.80 .3780 .6336 10.91 .0834 .7609
+DP-kernel $+$ MGS 19.41 .4349 .5261 16.47 .3854 .4621 14.87 .3050 .4806 20.31 .6097 .3676 11.57 .1617 .6643
+Real Scene Gradient Scaling Girl Heron Parterre Puppet Stair Average
+Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 9.41 .0744 .7648 11.73 .0565 .6825 14.10 .2013 .7586 12.72 .1821 .6405 13.87 .0276 .6817 13.26 .1797 .6732
+DP-kernel $+$ Naive [60] 9.35 .0906 .7656 11.64 .0645 .6473 13.97 .1795 .7477 12.64 .1865 .6280 14.10 .0280 .6827 13.20 .1789 .6680
+DP-kernel $+$ MGS 11.67 .2883 .5683 12.78 .1212 .5570 16.09 .2590 .5353 13.20 .2016 .6082 15.80 .1127 .5924 15.22 .2880 .5362
+TABLE XXVII: Quantitative comparison results for entire scenes from a 4-view setting that include MGS experiments on the world coordinate system. Color shading represents the best and second result.
+Synthetic Scene Gradient Scaling Factory Cozyroom Pool Tanabata Trolley Average
+Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 16.61 .3544 .5434 25.00 .7541 .1967 15.31 .1878 .8088 17.57 .4957 .3911 17.81 .4852 .4055 18.46 .4554 .4691
+DP-kernel $+$ Naive [60] 16.70 .3570 .5459 25.11 .7578 .2058 15.15 .1751 .7878 17.50 .4898 .3982 17.54 .4732 .4153 18.40 .4506 .4706
+DP-kernel $+$ MGS 16.71 .3612 .5380 24.48 .7441 .2072 16.63 .1934 .5907 16.90 .4590 .4038 17.96 .4887 .4105 18.54 .4493 .4300
+Real Scene Gradient Scaling Ball Basket Buick Coffee Decoration
+Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 22.39 .5417 .3903 19.18 .4771 .4105 14.51 .3651 .4723 21.86 .6689 .2854 11.54 .1274 .7296
+DP-kernel $+$ Naive [60] 22.84 .5624 .3842 18.70 .4611 .4302 16.33 .3849 .4132 21.05 .6320 .3154 11.17 .1222 .7454
+DP-kernel $+$ MGS 21.02 .4775 .4195 19.63 .4935 .4014 16.82 .4307 .4087 23.70 .7215 .2480 16.45 .3778 .4574
+Real Scene Gradient Scaling Girl Heron Parterre Puppet Stair Average
+Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 15.95 .5447 .3868 16.87 .3079 .3339 16.35 .2726 .5563 14.81 .2771 .4870 15.76 .0964 .5175 16.92 .3679 .4570
+DP-kernel $+$ Naive [60] 16.53 .5400 .3763 17.30 .3133 .3287 15.72 .2485 .5773 15.30 .2899 .4753 15.60 .0998 .4976 17.05 .3654 .4544
+DP-kernel $+$ MGS 16.31 .5390 .3694 17.04 .3202 .3195 16.70 .2837 .4821 15.01 .2914 .4659 16.39 .1255 .4470 17.91 .4061 .4019
+TABLE XXVIII: Quantitative comparison results for entire scenes from a 6-view setting that include MGS experiments on the world coordinate system. Color shading represents the best and second result.
+Synthetic Scene Gradient Scaling Factory Cozyroom Pool Tanabata Trolley Average
+Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 17.40 .3972 .4993 24.76 .7471 1878 16.10 .2033 .7592 18.35 .5266 .3951 18.60 .5370 .3628 19.04 .4822 .4408
+DP-kernel $+$ Naive [60] 17.36 .3929 .5039 24.27 .7370 .2001 16.33 .2201 .7567 18.67 .5289 .3888 18.42 .5363 .3628 19.01 .4830 .4425
+DP-kernel $+$ MGS 17.30 .3963 .4974 24.04 .7178 .1957 17.48 .2349 .5371 18.07 .5015 .3892 18.63 .5400 .3623 19.10 .4781 .3963
+Real Scene Gradient Scaling Ball Basket Buick Coffee Decoration
+Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 23.97 .6166 .3404 20.79 .5512 .3421 17.01 .4657 .3825 25.20 .7517 .2152 13.76 .2104 .6388
+DP-kernel $+$ Naive [60] 24.64 .6489 .3208 20.37 .5343 .3665 17.36 .4707 .3826 25.12 .7379 .2268 11.63 .1248 .7265
+DP-kernel $+$ MGS 23.86 .6111 .3449 20.91 .5549 .3511 17.78 .4768 .3802 26.10 .7655 .2119 16.60 .3919 .4509
+Real Scene Gradient Scaling Girl Heron Parterre Puppet Stair Average
+Kernel Type PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
+DP-kernel $\times$ 17.39 .5636 .3255 19.35 .4523 .2940 18.55 .4203 .4556 19.77 .5354 .3428 17.96 .1929 .4032 19.38 .4760 .3740
+DP-kernel $+$ Naive [60] 17.48 .5729 .3328 19.29 .4714 .2863 18.65 .4257 .4403 18.91 .4096 .3628 17.76 .1785 .3853 19.12 .4575 .3831
+DP-kernel $+$ MGS 18.08 .5934 .3184 18.37 .4190 .3011 18.91 .3998 .3986 19.63 .5319 .3443 17.69 .1798 .3682 19.79 .4923 .3470